### v2

# dark matter and baryon asymmetry of the Universe

Abstract
The electric dipole moment is examined in a three-loop neutrino mass model with dark matter originally proposed in Aoki et al [Phys. Rev. Lett. 102, 051805 (2009)]. The model contains a $CP$-violating phase in the Higgs potential which plays an important role in electroweak baryogenesis and is thus expected to explain the baryon asymmetry of the Universe simultaneously. However, such a $CP$-violating phase is severely constrained by the measurements of the electron electric dipole moment (eEDM), and a suppression mechanism for the eEDM is necessary to explain the observed baryon asymmetry while avoiding the constraint. In this paper, we examine neutrino mass, lepton-flavor-violating processes, dark matter, and the eEDM in the model. We show that the eEDM can be suppressed by destructive interference between the $CP$-violating phases in the Higgs sector and the dark sector with large $CP$-violating phases. We propose some benchmark scenarios including $O(1)$ $CP$-violating phases where tiny neutrino mass and dark matter can be explained while avoiding all current experimental and theoretical constraints. These $CP$-violating phases are expected to be large enough to generate the observed baryon asymmetry in the electroweak baryogenesis scenario.
preprint: OU-HET-1225
I Introduction
Although the standard model (SM) is consistent with various experimental results in particle physics and cosmology, we believe in the existence of physics beyond the SM because of its lack of ability to explain some phenomena such as neutrino oscillation Super-Kamiokande:1998kpq ; SNO:2001kpb , which implies tiny neutrino masses, the existence of dark matter (DM) Planck:2018vyg and baryon asymmetry of the Universe (BAU) Planck:2018vyg ; Fields:2019pfx . In recent decades, it has been one of the most important directions in particle physics to investigate new physics models to solve these problems.
From the theoretical perspective, the SM is an incomplete theory, i.e., there is no principle to determine the shape of the Higgs sector. Although the Higgs boson was discovered at LHC in 2012 ATLAS:2012yve , the entire structure of the Higgs sector has not been revealed, and there remains the possibility of an extended Higgs sector.
Given the above considerations, it would be natural to come up with a new physics model where an extended Higgs sector provides a mechanism to solve the above three problems. Various such mechanisms have been proposed so far. For example, loop-induced tiny neutrino masses via the quantum effect of additional scalar bosons Zee:1980ai , the scalar field DM as the weakly interacting massive particles (WIMP), and electroweak baryogenesis (EWBG) for the BAU Kuzmin:1985mm triggered by the strongly first-order electroweak phase transition (EWPT), which needs an extended Higgs sector Kajantie:1996mn .
Models for the loop-induced neutrino masses are called radiative seesaw models Zee:1980ai ; Cheng:1980qt ; Zee:1985id ; Krauss:2002px ; Ma:2006km ; Aoki:2008av ; Gustafsson:2012vj ; Aoki:2022bkg ; Nasri:2001ax ; Gu:2007ug ; Kanemura:2017haa ; Enomoto:2019mzl . In these models, the loop suppression factor naturally makes the size of neutrino masses small, and the mass of new particles can be at the TeV scale. In some radiative seesaw models, a new exact $Z_{2}$ symmetry is imposed to prohibit tree-level neutrino masses Krauss:2002px ; Ma:2006km ; Aoki:2008av ; Gustafsson:2012vj ; Aoki:2022bkg ; Nasri:2001ax ; Gu:2007ug ; Kanemura:2017haa ; Enomoto:2019mzl . This symmetry also plays a role in stabilizing the lightest $Z_{2}$-odd particle, which is a dark matter candidate. Thus, these models have an economical mechanism to explain the tiny neutrino mass and DM simultaneously.
In Ref. Aoki:2008av , a radiative seesaw model was proposed by Aoki, Kanemura and Seto (denoted by the AKS model in the following), where the BAU can also be explained in addition to tiny neutrino mass and DM by the Higgs sector extended by an additional Higgs doublet, a pair of charged scalar bosons and a real scalar boson. The latter two scalar bosons are dark sector particles that have an odd parity of an exact $Z_{2}$ symmetry. The extended Higgs sector can cause the strongly first-order EWPT and includes the $CP$-violating phase. Thus, baryon asymmetry can be generated via EWBG. In Ref. Aoki:2011zg , the Higgs potential of the AKS model and their theoretical constraints were studied in detail. However, in Refs. Aoki:2008av ; Aoki:2011zg , the $CP$-violating phases were neglected for simplicity. Thus, the BAU was not evaluated although Ref. Aoki:2008av numerically showed that the strongly first-order EWPT can be achieved in the model.
In Ref. Aoki:2022bkg , an extension of the AKS model was examined. It has a general Yukawa interaction between the SM fermions and the Higgs doublets, while those in the original AKS model are constrained by a softly broken $Z_{2}$ symmetry to avoid the flavor-changing neutral current (FCNC). By relaxing this symmetry, the model includes multiple $CP$-violating phases in not only the Higgs potential but also the Yukawa interactions, which make it possible to avoid the current severe constraint on the electron electric dipole moment (eEDM) ACME:2018yjb ; Roussy:2022cmp by the destructive interference between these $CP$-violating phases Kanemura:2020ibp . In Ref. Aoki:2022bkg , the authors performed a numerical analysis of the baryon number production via EWBG in the model and found one benchmark scenario where tiny neutrino masses, DM and the BAU can be simultaneously explained while avoiding all current experimental and theoretical constraints. However, Ref. Aoki:2022bkg assumes ad hoc flavor-alignment structure in the Yukawa interaction to restrict the FCNC Pich:2009sp , while the original AKS model naturally prohibits it at tree level by the softly broken $Z_{2}$ symmetry Glashow:1976nt . In addition, relaxing this symmetry causes too many model parameters.
In this paper, according to the facts mentioned above, we revisit the original AKS model to investigate the $CP$ violation. The most significant difficulty is avoiding the constraint from the eEDM measurements unless we assume that $CP$-violating phases are very small. We investigate the eEDM in the AKS model to see if it is possible to avoid the current severe bound while keeping a large $\mathcal{O}(1)$ $CP$-violating phase in the Higgs potential, which is necessary for successful EWBG Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg . The eEDM is induced by two kinds of Feynman diagrams at the leading order. One is generated by the additional Higgs bosons like in the two Higgs doublet models (THDMs) Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb , which depends on a $CP$-violating phase in the Higgs potential. The other one is generated by a new $CP$-violating Yukawa interaction between the charged leptons and the dark sector particles, which was neglected in Ref. Aoki:2022bkg . We consider a scenario where the eEDM is suppressed by the destructive interference between the Higgs sector contribution and the dark sector contribution to avoid the current severe constraint. This is a different suppression mechanism from that employed in Ref. Aoki:2022bkg , which was proposed in Ref. Kanemura:2020ibp and has been studied in Ref. Kanemura:2021atq ; Enomoto:2021dkl ; Kanemura:2023juv ; Kanemura:2023jbz . It is also different from other conventional suppression mechanisms in the THDMs Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb .
We also examine the neutrino mass, lepton-flavor-violating (LFV) processes, and DM in the model and show two benchmark scenarios where tiny neutrino masses and DM can be simultaneously explained while avoiding all current experimental and theoretical constraints. In the two scenarios, DM is the real scalar boson, but with different masses. In the first scenario, the mass is about half of the 125 GeV Higgs boson, and a pair of DM mainly annihilates into the SM fermions via an almost on-shell Higgs boson. On the other hand, in the second scenario, DM has heavier mass, and a pair of them can annihilate into the additional Higgs bosons. In both scenarios, the size of the eEDM is smaller than the current limit while keeping an $\mathcal{O}(1)$ $CP$-violating phase in the Higgs potential.
This paper is organized as follows. In the following section (Sec. II), we define the model and show the lagrangian. Theoretical and experimental constraints on the model are also shown there. In Sec. III, we discuss some phenomenologies in the model: neutrino mass (Sec. III.1), lepton flavor violation (III.2) and DM (III.3). In Sec. IV, the eEDM in the model is investigated. Sec. V is assigned to discussions about numerical evaluations of the neutrino mass, the LFV processes, DM and the eEDM. We show two benchmark scenarios that can explain tiny neutrino masses and DM simultaneously while avoiding all the experimental and theoretical constraints. Some comments about the benchmark scenarios and a conclusion are presented in Sec. VI.
II The model
II.1 Lagrangian
Here, we show the lagrangian of the AKS model with $CP$-violation. The notation follows Ref. Aoki:2008av . The model includes new discrete symmetries; an exact $Z_{2}$ symmetry and a softly broken $Z_{2}$ symmetry, which are denoted by $Z_{2}$ and $\tilde{Z}_{2}$, respectively in the following. The gauge symmetry is the same as in the SM. The matter fields and their quantum numbers are shown in Table 1. There are three kinds of $Z_{2}$-odd fields: a real scalar $\eta$, a pair of charged scalars $S^{\pm}$, and three right-handed fermions $N_{R}^{\alpha}$ ($\alpha=1,2,3$). $N_{R}^{\alpha}$ and $\eta$ are gauge-singlet, while $S^{\pm}$ has a hypercharge ±$\pm 1$, respectively. $N^{\alpha}_{R}$ constitute Majorana fermions $N^{\alpha}\equiv N_{R}^{\alpha}+(N_{R}^{\alpha})^{c}$ with the mass
$\mathcal{L}=$ $-\frac{m_{N^{\alpha}}}{2}\overline{(N_{R}^{\alpha})^{c}}N_{R}^{\alpha}+\mathrm{h.c.}=-\frac{m_{N^{\alpha}}}{2}(N^{\alpha})^{\mathrm{T}}CN^{\alpha},$ (1)
where $C$ is the charge conjugate matrix. The SM fermions have $Z_{2}$-even parities, and we do not have a new $Z_{2}$-even fermion. On the other hand, $Z_{2}$-even scalar sector is an extension of the SM and includes two Higgs doublets $\phi_{1}$ and $\phi_{2}$, which have an even and odd parity of $\tilde{Z}_{2}$, respectively to avoid the FCNCs Glashow:1976nt .
$Q_{L}^{\prime i}$ $u_{R}^{\prime i}$ $d_{R}^{\prime i}$ $L_{L}^{\prime i}$ $\ell_{R}^{\prime i}$ $\phi_{1}$ $\phi_{2}$ $N_{R}^{\alpha}$ $S^{+}$ $\eta$
Spin $1/2$ $1/2$ $1/2$ $1/2$ $1/2$ 00 00 $1/2$ 00 00
$SU(3)_{C}$ 3 3 3 1 1 1 1 1 1 1
$SU(2)_{L}$ 2 1 1 2 1 2 2 1 1 1
$U(1)_{Y}$ $1/6$ $2/3$ $-1/3$ $-1/2$ $-1$ $1/2$ $1/2$ 00 $1$ 00
$Z_{2}$ $+$ --
$\tilde{Z}_{2}$ $+$ -- -- $+$ $+$ $+$ -- $+$ $+$ $+$
Table 1: Matter fields in the model and their quantum numbers. $Z_{2}$ and $\tilde{Z}_{2}$ are exact and softly-broken $Z_{2}$ symmetries, respectively.
The Higgs potential $V$ is given by
$V=$ $\ V_{\mathrm{THDM}}+\mu_{S}^{2}|S^{+}|^{2}+\frac{\mu_{\eta}^{2}}{2}\eta^{2}+\frac{\lambda_{S}}{4}|S^{+}|^{4}+\frac{\lambda_{\eta}}{4!}\eta^{4}+\frac{\xi}{2}|S^{+}|^{2}\eta^{2}$
$+\sum_{a=1}^{2}|\phi_{a}|^{2}\Bigl{(}\rho_{a}|S^{+}|^{2}+\frac{\sigma_{a}}{2}\eta^{2}\Bigr{)}+\Bigl{(}2\kappa\tilde{\phi}_{1}^{\dagger}\phi_{2}S^{-}\eta+\mathrm{h.c.}\Bigr{)}.$ (2)
$V_{\mathrm{THDM}}$ is the potential in the $CP$-violating THDM;
$V_{\mathrm{THDM}}=\sum_{a=1}^{2}\biggl{(}-\mu_{a}^{2}|\phi_{a}|^{2}+\frac{\lambda_{a}}{2}|\phi_{a}|^{4}\biggr{)}-\bigl{(}\mu_{12}^{2}\phi_{1}^{\dagger}\phi_{2}+\mathrm{h.c.}\bigr{)}$
$+\lambda_{3}|\phi_{1}|^{2}|\phi_{2}|^{2}+\lambda_{4}|\phi_{1}^{\dagger}\phi_{2}|^{2}+\biggl{(}\frac{\lambda_{5}}{2}(\phi_{1}^{\dagger}\phi_{2})^{2}+\mathrm{h.c.}\biggr{)}.$ (3)
where $\tilde{\phi}_{1}=i\sigma_{2}\phi_{1}^{\ast}$. In general, $\mu_{12}^{2}$ and $\lambda_{5}$, and $\kappa$ are complex. The rephasing of $S^{\pm}$ can make $\kappa$ real without loss of generality. On the other hand, the phase of $\mu_{12}^{2}$ and $\lambda_{5}$ remains as $CP$-violating phases as discussed below.
The elements of the Higgs doublets are defined as
$\phi_{i}=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}w^{+}_{i}\\ v_{i}e^{i\xi_{i}}+h_{i}+ia_{i}\\ \end{pmatrix}\quad i=1,2,$ (4)
where $v_{1}$ and $v_{2}$ are positive constant. In general, the vacuum expectation values (VEVs) of the doublets are complex, i.e., the phase $\xi_{i}$ are not zero or $\pi$. However, only the relative phase $\xi_{21}=\xi_{2}-\xi_{1}$ are physical because of the $U(1)_{Y}$ symmetry. We thus take $\xi_{1}=0$ and $\xi_{2}=\xi_{21}$.
We have three complex phases $\theta_{\mu}=\mathrm{arg}[\mu_{12}^{2}]$, $\theta_{5}=\mathrm{arg}[\lambda_{5}]$, and $\xi_{21}$ in the Higgs potential. One of them can be real by rephasing $\phi_{2}$, and there remain two $CP$-violating phases. We choose $\xi_{21}=0$ in the following.
The stationary condition at the vacuum requires the following equations to be satisfied;
[…]
where $\lambda_{345}^{+}=\lambda_{3}+\lambda_{4}+\lambda_{5}^{R}$, and the superscripts $R$ and $I$ represents the real and imaginary part of each quantity. We note that $\theta_{\mu}$ and $\theta_{5}$ are related by the third equation. Therefore, there is only one independent $CP$-violating phase in the Higgs sector. We choose $\theta_{5}$ as an input parameter. Then, $\theta_{\mu}$ is determined by the third equation. In addition, we note that dimension-two paramters $\mu_{1}^{2}$, $\mu_{2}^{2}$, and $\mu_{12}^{2I}$ are replaced by other paramters by the above equations. Thus, $\mu_{12}^{2R}$ is a unique dimensionful free parameter in $V_{\mathrm{THDM}}$.
To investigate mass eigenstates of the Higgs bosons, it is convenient to change the basis of the doublet from $\tilde{Z}_{2}$-parity eigenstate basis ($\phi_{1}$, $\phi_{2}$) to the Higgs basis ($\Phi_{1}$, $\Phi_{2}$), where only $\Phi_{1}$ obtains the real VEV Davidson:2005cw . This transformation is achieved by the rotation
$\begin{pmatrix}\Phi_{1}\\ \Phi_{2}\\ \end{pmatrix}=\begin{pmatrix}\cos\beta&\sin\beta\\ -\sin\beta&\cos\beta\\ \end{pmatrix}\begin{pmatrix}\phi_{1}\\ \phi_{2}\\ \end{pmatrix},$ (6)
where $\beta$ satisfies $\tan\beta=v_{2}/v_{1}$.
$\Phi_{1}$ and $\Phi_{2}$ are given by
$\Phi_{1}=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}G^{+}\\ v+h_{1}^{\prime}+G^{0}\\ \end{pmatrix},\quad\Phi_{2}=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}H^{+}\\ h_{2}^{\prime}+ih_{3}^{\prime}\\ \end{pmatrix},$ (7)
where $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\simeq 246~{}\mathrm{GeV}$. Then, $G^{+}$ and $G^{0}$ are Nambu-Goldstone (NG) bosons and are absorbed into the longitudinal model of $W^{\pm}$ and $Z$ bosons, respectively. $H^{\pm}$ is a physical charged scalar boson, and its mass $m_{H^{\pm}}$ is given by
$m_{H^{\pm}}^{2}=M^{2}-\frac{v^{2}}{2}(\lambda_{4}+\lambda_{5}^{R}),$ (8)
where $M^{2}=\mu_{12}^{2R}/(s_{\beta}c_{\beta})$. Neutral scalar bosons $h_{1}^{\prime}$, $h_{2}^{\prime}$, and $h_{3}^{\prime}$ are not mass eigenstates, and their mass matrix is given by
$M_{h}^{2}=v^{2}\begin{pmatrix}\lambda_{1}c_{\beta}^{4}+\lambda_{2}s_{\beta}^{4}+2\lambda_{345}^{+}s_{\beta}^{2}c_{\beta}^{2}&\bigl{\{}\lambda_{2}s_{\beta}^{2}-\lambda_{1}c_{\beta}^{2}+\lambda_{345}^{+}c_{2\beta}\bigr{\}}s_{\beta}c_{\beta}&-\lambda_{5}^{I}s_{\beta}c_{\beta}\\[10.0pt] \bigl{\{}\lambda_{2}s_{\beta}^{2}-\lambda_{1}c_{\beta}^{2}+\lambda_{345}^{+}c_{2\beta}\bigr{\}}s_{\beta}c_{\beta}&\displaystyle{\frac{M^{2}}{v^{2}}+(\lambda_{1}+\lambda_{2}-2\lambda_{345}^{+})s_{\beta}^{2}c_{\beta}^{2}}&\displaystyle{-\frac{1}{2}\lambda_{5}^{I}c_{2\beta}}\\[10.0pt] -\lambda_{5}^{I}s_{\beta}c_{\beta}&\displaystyle{-\frac{1}{2}\lambda_{5}^{I}c_{2\beta}}&\displaystyle{\frac{M^{2}}{v^{2}}-\lambda_{5}^{R}}\\ \end{pmatrix},$ (9)
where $s_{\beta}=\sin\beta$, $c_{\beta}=\cos\beta$ and $c_{2\beta}=\cos 2\beta$. According to Refs. Kanemura:2015ska , we use the following parametrization;
$(M_{h}^{2})_{11}=\tilde{m}_{h}^{2}s_{\beta-\tilde{\alpha}}^{2}+\tilde{m}_{H}^{2}c^{2}_{\beta-\tilde{\alpha}},$ (10)
$(M_{h}^{2})_{12}=(M_{h}^{2})_{21}=(\tilde{m}_{h}^{2}-\tilde{m}_{H}^{2})s_{\beta-\tilde{\alpha}}c_{\beta-\tilde{\alpha}},$ (11)
$(M_{h}^{2})_{22}=\tilde{m}_{h}^{2}c_{\beta-\tilde{\alpha}}^{2}+\tilde{m}_{H}^{2}s^{2}_{\beta-\tilde{\alpha}},$ (12)
$(M_{h}^{2})_{33}=\tilde{m}_{A}^{2},$ (13)
where $s_{\beta-\tilde{\alpha}}=\sin(\beta-\tilde{\alpha})$ and $c_{\beta-\tilde{\alpha}}=\cos(\beta-\tilde{\alpha})$. We note that $\tilde{m}_{h}$, $\tilde{m}_{H}$, and $\tilde{m}_{A}$ are the mass eigenvalues in the $CP$-conserving limit ($\lambda_{5}^{I}=0$), where $h_{3}^{\prime}$ is the $CP$-odd mass eigenstate, and $h_{1}^{\prime}$ and $h_{2}^{\prime}$ are mixed by the mixing angle $\beta-\tilde{\alpha}$.
On the other hand, in the $CP$-violating case, $\lambda_{5}^{I}$ generates additional mixings, and $\tilde{m}_{h}$, $\tilde{m}_{H}$, and $\tilde{m}_{A}$ are no longer mass eigenvalues. We define the mass eigenvalues and eigenstates as follows;
$\mathrm{diag}(m_{H_{1}}^{2},m_{H_{2}}^{2},m_{H_{3}}^{2})=R^{\mathrm{T}}M_{h}^{2}R,\quad(H_{1},H_{2},H_{3})=(h_{1}^{\prime},h_{2}^{\prime},h_{3}^{\prime})R,$ (14)
where the $R$ is an appropriate orthogonal matrix. Since it is difficult to solve the matrix $R$ analytically, we use $\tilde{m}_{H}$ and $\tilde{m}_{A}$ as input parameters, not $m_{H_{2}}$ and $m_{H_{3}}$. On the other hand, $\tilde{m}_{h}$ is determined so that $m_{H_{1}}=125~{}\mathrm{GeV}$, and $H_{1}$ is the SM Higgs boson. Then, the mass matrix $M_{h}^{2}$ includes five free parameters: $\tilde{m}_{H}$, $\tilde{m}_{A}$, $\tan\beta$, $s_{\beta-\tilde{\alpha}}$, and $\lambda_{5}^{I}$.
Once the mass matrix is fixed, we can find the mass eigenvalues and the mixing matrix $R$. In addition, we can determine scalar couplings as follows Kanemura:2015ska ;
$\lambda_{1}v^{2}=(M_{h}^{2})_{11}+\bigl{(}(M_{h}^{2})_{22}-M^{2}\bigr{)}t_{\beta}^{2}-2(M_{h}^{2})_{12}t_{\beta},$ (15)
$\lambda_{2}v^{2}=(M_{h}^{2})_{11}+\bigl{(}(M_{h}^{2})_{22}-M^{2}\bigr{)}t_{\beta}^{-2}+2(M_{h}^{2})_{12}t_{\beta}^{-1},$ (16)
$\lambda_{3}v^{2}=(M_{h}^{2})_{11}-\bigl{(}(M_{h}^{2})_{22}+M^{2}\bigr{)}\cot 2\beta+2m_{H^{\pm}}^{2},$ (17)
$\lambda_{4}v^{2}=M^{2}+(M_{h}^{2})_{33}-2m_{H^{\pm}}^{2},$ (18)
$\lambda_{5}^{R}v^{2}=M^{2}-(M_{h}^{2})_{33},$ (19)
where $t_{\beta}=\tan\beta$. Consequently, $V_{\mathrm{THDM}}$ has the following free paraemeters;
$M,\quad\tilde{m}_{H},\quad\tilde{m}_{A},\quad m_{H^{\pm}},\quad\tan\beta,\quad\sin(\beta-\tilde{\alpha}),\quad\theta_{5}.$ (20)
Next, we consider the $Z_{2}$-odd scalar bosons. The masses of $S^{\pm}$ and $\eta$ are given by
$m_{S}^{2}=\mu_{S}^{2}+\frac{v^{2}}{2}(\rho_{1}c_{\beta}^{2}+\rho_{2}s_{\beta}^{2}),$ (21)
$m_{\eta}^{2}=\mu_{\eta}^{2}+\frac{v^{2}}{2}(\sigma_{1}c_{\beta}^{2}+\sigma_{2}s_{\beta}^{2}).$ (22)
We use $m_{S}$ and $m_{\eta}$ as input parameters instead of $\mu_{S}$ and $\mu_{\eta}$. Thus, the remaining part of the scalar potential is determined by
$m_{S},\quad m_{\eta},\quad\lambda_{S},\quad\lambda_{\eta},\quad\xi,\quad\rho_{a},\quad\sigma_{a},\quad\kappa,$ (23)
where $a=1,2$. As a result, the Higgs potential has five dimensionful parameters, two mixing angles, six dimensionless couplings and one $CP$-violating phase.
Next, we consider the Yukawa interactions. Because of the $\tilde{Z}_{2}$ parity of the SM fermions shown in Table 1, the Yukawa interactions between the Higgs doublets and the SM fermions are given by
$\mathcal{L}=$ $-(y_{u})_{ij}\overline{Q_{L}^{\prime i}}\tilde{\phi}_{2}u_{R}{\prime j}-(y_{d})_{ij}\overline{Q_{L}^{\prime i}}\phi_{2}d_{R}{\prime j}-(y_{\ell})_{ij}\overline{L_{L}^{\prime i}}\phi_{1}\ell_{R}{\prime j}+\mathrm{h.c.},$ (24)
which is the same as in the Type-X THDM Logan:2009uf ; Su:2009fz ; Grossman:1994jb ; Aoki:2009ha ; Barger:1989fj . After the bi-unitary transformations of the SM fermions, the Yukawa interactions are given by
[…]
$+\frac{m_{\ell^{i}}}{v}\overline{\ell^{i}}\Bigl{(}H_{1}R_{1i}+H_{2}R_{2i}t_{\beta}+iH_{3}R_{3i}t_{\beta}\gamma_{5}\Bigr{)}\ell^{i}\biggr{\}}$
[…]
The model has an additional Yukawa coupling among $S^{\pm}$, $N^{\alpha}_{R}$, and $\ell_{R}^{i}$ as follows;
$\mathcal{L}=-\sum_{i,\alpha}h_{i}^{\alpha}\overline{(N_{R}^{\alpha})^{c}}\ell_{R}^{i}S^{+}+\mathrm{h.c.}=-\sum_{i,\alpha}h_{i}^{\alpha}\overline{(N^{\alpha})^{c}}P_{R}\ell^{i}S^{+}+\mathrm{h.c.}.$ (26)
We note that the coupling $h_{i}^{\alpha}$ is defined for the mass eigenstates of the charged leptons (without ). The matrix $h$ includes nine complex parameres. Three of them can be real by rephasing the lepton fields. This degree of freedom is the same as those to reduce the three unphysical phases in the (Pontecorvo-Maki-Nakagawa-Sakata) matrix $U$ ref: Pontecorvo ; Maki:1962mu . We have a detailed discussion about in Sec. III.1.
II.2 Theoretical constraints
Here, we discuss theoretical constraints on the Higgs potential, i.e., the vacuum stability, the triviality bound, and the perturbative unitarity. To ensure that the local minimum is the stable vacuum of theory, a necessary condition is normally imposed such that the Higgs potential does not fall into negative infinity with a large field configuration along any direction of the scalar fields. It requires some combinations of quartic scalar couplings to be positive. According to Ref. Aoki:2011zg , we employ the following conditions;
$\lambda_{1}>0,\quad\lambda_{2}>0,\quad\lambda_{S}>0,\quad\lambda_{\eta}>0,$ (27)
$\sqrt{\lambda_{1}\lambda_{2}}+\lambda_{3}+\mathrm{min}(0,\lambda_{4}-|\lambda_{5}|)>0,$ (28)
$\sqrt{\lambda_{1}\lambda_{S}/2}+\rho_{1}>0,\quad\sqrt{\lambda_{1}\lambda_{\eta}/3}+\sigma_{1}>0,\sqrt{\lambda_{2}\lambda_{S}/2}+\rho_{2}>0,$ (29)
$\sqrt{\lambda_{2}\lambda_{\eta}/3}>0,\quad\sqrt{\lambda_{S}\lambda_{\eta}/6}+\xi>0,$ (30)
$0\displaystyle 2\lambda_{1}+2\lambda_{2}+4\lambda_{3}+4\rho_{1}+4\rho_{2}+\lambda_{S}+4\sigma_{1}+4\sigma_{2}+\frac{2}{3}\lambda_{\eta}+4\xi-16\sqrt{2}|\kappa|>$ 0 . (31)
In Eq. (28), we use $|\lambda_{5}|$, not $\lambda_{5}^{R}$ shown in Ref. Aoki:2011zg , to include the effect of the $CP$-violation in $\lambda_{5}$ according to discussion in the THDMs Deshpande:1977rw ; Klimenko:1984qx ; Nie:1998yn ; Ferreira:2004yd .
The triviality bound requires that all the running couplings do not infinitely blow up nor fall down below the cut-off scale $\Lambda$. Demanding $\Lambda$ be higher than a certain value, we find constraints on the scalar couplings. Ref. Aoki:2011zg studies the triviality bound in the AKS model by using the mass-independent renormalization group equation (RGE), where $\Lambda$ strongly depends on a scale from which the new physics effect is included in the running (the threshold effect).
On the other hand, Ref. Kanemura:2023wap used the mass-dependent renormalization group equation to study the triviality bound in some new physics models with extended Higgs sectors. Then, the new physics effect naturally decouples in the running at low energies without setting an artificial threshold. They also compare the result with that in the mass-independent RGE and study how the threshold effect changes $\Lambda$. According to Refs. Aoki:2011zg ; Kanemura:2023wap , we require the size of all the scalar couplings to be less than three to make $\Lambda$ higher than $10~{}\mathrm{TeV}$.
Demanding the perturbative unitarity constrains the quartic scalar couplings so that the s-wave amplitude $a_{0}$ satisfies $|a_{0}|<c$ for all two-to-two-body scalar boson scattering Lee:1977eg , where $c=1$ Lee:1977eg or $1/2$ Luscher:1988gc . It leads to an upper bound for the size of the scalar couplings, which implies that masses of new particles are not far from their invariant mass parameters, i.e., $M^{2}$, $\mu_{S}^{2}$, and $\mu_{\eta}^{2}$ Kanemura:1993hm ; Ginzburg:2005dt ; Akeroyd:2000wc ; Kanemura:2015ska . We use the result in Ref. Kanemura:2015ska and requires the following conditions for the quartic couplings in $V_{\mathrm{THDM}}$;
$3(\lambda_{1}+\lambda_{2})\pm\sqrt{9(\lambda_{1}-\lambda_{2})^{2}+4(2\lambda_{3}+\lambda_{4})^{2}}<16\pi,$ (32)
$2\lambda_{3}+2\lambda_{4}\pm 3\lambda_{5}^{R}<16\pi,$ (33)
$\lambda_{1}+\lambda_{2}\pm\sqrt{(\lambda_{1}-\lambda_{2})^{2}+4\lambda_{4}^{2}}<16\pi,$ (34)
$2\lambda_{3}\pm 2|\lambda_{5}|<16\pi,$ (35)
$\lambda_{1}+\lambda_{2}\pm\sqrt{(\lambda_{1}-\lambda_{2})^{2}+4|\lambda_{5}|^{2}}<16\pi,$ (36)
$\lambda_{4}+\lambda_{3}<16\pi,$ (37)
where we employ $c=1/2$. For other coupling in the Higgs potential, we simply require them to be less than three to satisfy the perturbative unitarity.
II.3 Experimental constraints
In this subsection, we describe experimental constraints on the new particles. For simplicity, we neglect the effect of the $CP$-violating couplings. We assume that $m_{N^{\alpha}}=\mathcal{O}(1)~{}\mathrm{TeV}$, and the masses of the other particles are $\mathcal{O}(100)~{}\mathrm{GeV}$. Then, $\eta$ is the dark matter particle. We use the theoretical results in Refs. Aiko:2020ksl ; Arbey:2017gmh and Enomoto:2015wbn ; Arbey:2017gmh ; Haller:2018nnx to discuss constraints on the additional Higgs bosons from the collider experiments and flavor experiments, respectively.
$H^{\pm}$ has the Type-X (lepton specific) Yukawa coupling with the SM fermions Logan:2009uf ; Su:2009fz ; Grossman:1994jb ; Aoki:2009ha ; Barger:1989fj . The LEP result gives the lower limit $m_{H^{\pm}}\gtrsim 80~{}\mathrm{GeV}$, which is almost independent of $\tan\beta$ ALEPH:2013htx . At LHC, the main production channel for light $H^{\pm}$ is the top quark decay in the case that $m_{H^{\pm}}<m_{t}-m_{b}\simeq 170~{}\mathrm{GeV}$. The strong constraint comes from $H^{\pm}\to\tau\nu$ search, and it is $\tan\beta\gtrsim 10$ Arbey:2017gmh . If $H^{\pm}$ is larger than $170~{}\mathrm{GeV}$, it is produced by the associated production $gg\to tbH^{\pm}$, and $H^{\pm}$ predominantly decays into $\tau\nu$ or $tb$. The $tb$ search gives a stronger constraint; $\tan\beta\gtrsim 2$ (1.5) for $m_{H^{\pm}}=200$ (400) GeV Aiko:2020ksl .
$H^{\pm}$ is also constrained by flavor observables Haller:2018nnx ; Enomoto:2015wbn ; Arbey:2017gmh such as $B\to X_{s}\gamma$ HFLAV:2019otj , $B_{s}$($B_{d}$) $\to\mu\mu$ ATLAS:2018cur . In the current case, the strongest constraint comes from $B_{d}\to\mu\mu$ Haller:2018nnx ; Enomoto:2015wbn ; Arbey:2017gmh , and it is $\tan\beta\gtrsim 3$ and $2$ for $m_{H^{\pm}}=100$ and $400~{}\mathrm{GeV}$, respectively. Thus, the flavor experiment gives a stronger constraint than the collider experiments for $H^{\pm}$ heavier than $170~{}\mathrm{GeV}$ Arbey:2017gmh .
$H_{2}$ and $H_{3}$ are predominantly produced via the gluon fusion $gg\to H_{2,3}$ at LHC because we consider the type-X Yukawa interaction Aiko:2020ksl . The constraint strongly depends on their mixing with the SM Higgs boson $H_{1}$. In the case without mixing (the alignment case), they are constrained by the searches for $H_{2,3}\to\tau^{+}\tau^{-}$ and $H_{2,3}\to t\bar{t}$. If they mix with $H_{1}$, in addition to these two processes, constraints from $H_{2}\to ZZ$, $H_{2}\to H_{1}H_{1}$ and $H_{3}\to ZH_{1}$ are important. All of these constraints are relaxed by considering large $\tan\beta$ because the production cross section is proportional to $\cot\beta$. Here, we employ $\tan\beta\gtrsim 15$ as a constraint for $H_{2,3}$ with the mass of a few hundred GeV.
[…]
$N^{\alpha}$ couple only the charged leptons. Thus, the production at hadron colliders is suppressed. At $e^{+}e^{-}$ colliders, a pair of them can be produced via the t–channel diagram mediated by $S^{\pm}$. In the following, we are interested in $N^{\alpha}$ with the mass $m_{N^{\alpha}}\simeq\mathcal{O}(1)~{}\mathrm{TeV}$. Then, $N^{\alpha}$ are not constrained by the collider experiments so far.
The mass of $S^{\pm}$ and $N^{\alpha}$ and the size of the Yukawa coupling $h_{i}^{\alpha}$ are strongly constrained by the current measurements of LFV processes. We discuss this constraint in detail in Sec. V.
III Phenomenology
In this section, we discuss the phenomenology of the model, neutrino mass, lepton flavor violating decays, and dark matter. Although these are already studied in previous works Aoki:2008av ; Aoki:2011zg ; Aoki:2022bkg , there are several updates about the parametrization of $h_{i}^{\alpha}$ and the dark matter phenomenology.
III.1 Neutrino mass
Figure 1: Feynman diagram for neutrino mass
In the model, Majorana masses of neutrinos are generated by the three-loop diagram in Fig. 1. The mass matrix is given by
$(M_{\nu})_{ij}=\frac{(\kappa\tan\beta)^{2}m_{\ell^{i}}m_{\ell^{j}}}{(16\pi^{2})^{3}}h_{i}^{\alpha}h_{j}^{\alpha}m_{N^{\alpha}}\Bigl{(}F_{1\alpha}+F_{2\alpha}\Bigr{)},$ (38)
where $F_{1\alpha}$ and $F_{2\alpha}$ are loop functions corresponding to the left and right diagrams in Fig. 1, respectively. They are given by
$F_{n\alpha}=\int_{0}^{1}\tilde{\mathrm{d}}^{4}x\int_{0}^{\infty}\mathrm{d}u\int_{0}^{\infty}\mathrm{d}v\frac{8\sqrt{uv}\tilde{F}(a_{n},b_{n})}{(u+m_{H^{\pm}}^{2})(v+m_{H^{\pm}}^{2})},$ (39)
where $n=1,2$, and
$\int\mathrm{\tilde{d}}^{4}x=\int_{0}^{1}\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}z\,\mathrm{d}\omega\ \delta(x+y+z+\omega-1).$ (40)
The function $\tilde{F}$ and the variables $a_{n}$ and $b_{n}$ are defined as
$\tilde{F}(x,y)=\frac{1}{y^{3}}\Bigl{(}\sqrt{x^{2}-y^{2}}+\frac{x^{2}}{\sqrt{x^{2}-y^{2}}}-2x\Bigr{)},$ (41)
$a_{1}=(y+z)m_{S}^{2}+xm_{N^{\alpha}}^{2}+\omega m_{\eta}^{2}+z(1-z)u+y(1-y)v,$ (42)
$a_{2}=(y+z)m_{S}^{2}+xm_{N^{\alpha}}^{2}+\omega m_{\eta}^{2}+(y+\omega)(x+z)u+(x+y)(z+\omega)v,$ (43)
$b_{1}=2yz\sqrt{uv},$ (44)
$b_{2}=2(yz-x\omega)\sqrt{uv}.$ (45)
See Ref. Aoki:2022bkg for the derivation of the formulas.
Next, we consider the parametrization of $h_{i}^{\alpha}$ to explain the neutrino oscillation data. It is also convenient to discuss the $CP$-violating phases in $h_{i}^{\alpha}$. The formula in Eq. (38) can be written by the following simple matrix form;
$M_{\nu}=Y^{\mathrm{T}}\Lambda^{-1}Y,$ (46)
where $Y$ and $\Lambda$ are $3\times 3$ matrices defined as
$Y_{\alpha\ell}=\frac{|\kappa\tan\beta|}{(16\pi^{2})^{3/2}}h^{\alpha}_{i}m_{\ell^{i}},\quad\Lambda_{\alpha\beta}=\Lambda_{\alpha}\delta_{\alpha\beta},$ (47)
with
$\Lambda_{\alpha}=\bigl{\{}m_{N^{\alpha}}(F_{1\alpha}+F_{2\alpha})\bigr{\}}^{-1}.$ (48)
We note that $\Lambda_{\alpha}$ has a mass dimension one and is a typical new physics scale for the neutrino mass. $M_{\nu}$ is diagonalized by the PMNS matrix $U$,
$D\equiv\mathrm{diag}(m_{\nu^{1}},m_{\nu^{2}},m_{\nu^{3}})=U^{\mathrm{T}}Y^{\mathrm{T}}\Lambda^{-1}YU.$ (49)
Since the structure of the neutrino mass matrix in Eq. (46) is the same as that in the type-I seesaw mechanism ref: Type-I seesaw , we can use the Casas-Ibarra parametrization Casas:2001sr to study parameter regions to explain the neutrino oscillation data. Then, the matrix $Y$ is given by $Y=\sqrt{|\Lambda|}O\sqrt{D}U^{\dagger}$, where the matrix $O$ is a complex orthogonal matrix whose elements are independent of the neutrino oscillation data, $\sqrt{|\Lambda|}=\mathrm{diag}(\sqrt{|\lambda_{1}|},\sqrt{|\lambda_{2}|},\sqrt{|\lambda_{3}|})$ and $\sqrt{D}=\mathrm{diag}(\sqrt{m_{\nu^{1}}},\sqrt{m_{\nu^{2}}},\sqrt{m_{\nu^{3}}})$. Consequently, the Yukawa matrix $h$ has to be the following form to explain the neutrino oscillation data;
$h=\frac{(16\pi^{2})^{3/2}}{|\kappa\tan\beta|}\sqrt{|\Lambda|}O\sqrt{D}U^{\dagger}M_{L}^{-1},$ (50)
where $M_{L}=\mathrm{diag}(m_{e},m_{\mu},m_{\tau})$.
The matrix $O$ consists of three complex angles $z_{1}$, $z_{2}$, and $z_{3}$ as
$O=\begin{pmatrix}1&0&0\\ 0&c_{z_{1}}&s_{z_{1}}\\ 0&-s_{z_{1}}&c_{z_{1}}\\ \end{pmatrix}\begin{pmatrix}c_{z_{2}}&0&s_{z_{2}}\\ 0&1&0\\ -s_{z_{2}}&0&c_{z_{2}}\\ \end{pmatrix}\begin{pmatrix}c_{z_{3}}&s_{z_{3}}&0\\ -s_{z_{3}}&c_{z_{3}}&0\\ 0&0&1\\ \end{pmatrix},$ (51)
where $c_{X}=\cos X$ and $s_{X}=\sin X$ for any $X$. In this parametrization, the imaginary parts of $z_{i}$ are unbounded. Instead, we employ the following formula with six real angles;
$O=$ $\frac{1}{c_{\psi_{1}}c_{\psi_{2}}c_{\psi_{3}}}\begin{pmatrix}1&0&0\\ 0&c_{\gamma_{1}}&s_{\gamma_{1}}\\ 0&-s_{\gamma_{1}}&c_{\gamma_{1}}\\ \end{pmatrix}\begin{pmatrix}c_{\psi_{1}}&0&0\\ 0&1&is_{\psi_{1}}\\ 0&-is_{\psi_{1}}&1\\ \end{pmatrix}\begin{pmatrix}c_{\gamma_{2}}&0&s_{\gamma_{2}}\\ 0&1&0\\ -s_{\gamma_{2}}&0&c_{\gamma_{2}}\\ \end{pmatrix}$
×$\times\begin{pmatrix}1&0&is_{\psi_{2}}\\ 0&c_{\psi_{2}}&0\\ -is_{\psi_{2}}&0&1\\ \end{pmatrix}\begin{pmatrix}c_{\gamma_{3}}&s_{\gamma_{3}}&0\\ -s_{\gamma_{3}}&c_{\gamma_{3}}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}1&is_{\psi_{3}}&0\\ -is_{\psi_{3}}&1&0\\ 0&0&c_{\psi_{3}}\\ \end{pmatrix},$ (52)
where we used $\gamma_{i}=\mathrm{Re}[z_{i}]$, $\cosh(\mathrm{Im}[z_{i}])=\cos^{-1}\psi_{i}$, and $\sinh(\mathrm{Im}[z_{i}])=\tan\psi_{i}$ for $i=1$, 2, 3. Then, $\gamma_{i}$ and $\psi_{i}$ are bounded in the regions $[0,2\pi)$ and $[-\pi/2,\pi/2]$, respectively. In the limit of $\psi_{i}\to 0$, the matrix $O$ is a real orthogonal matrix.
We revisit the $CP$-violating phases in $h_{i}^{\alpha}$ by using Eq. (50). As explained in Sec. II.1, the matrix $h$ generally includes nine complex phases, but three of them can be zero by rephasing the lepton fields. By using Eq. (50), we can understand the fact that this degree of freedom is the same as that to reduce unphysical phases from the PMNS matrix. The remaining $CP$-violating phases in $U_{ia}$ are the Dirac $CP$ phase $\delta$ and two Majorana $CP$ phases $\alpha_{1}$ and $\alpha_{2}$. As a result, there are six $CP$-violating phases in $h_{i}^{\alpha}$; $\delta$, $\alpha_{1}$ and $\alpha_{2}$ in $U_{ia}$ and $\psi_{1}$, $\psi_{2}$ and $\psi_{3}$ in $O_{a\alpha}$.
III.2 Lepton flavor violation
In the model, the lepton flavor violating decay $\ell^{i}\to\ell^{j}\gamma$ occurs at one-loop level via the Yukawa coupling $h_{i}^{\alpha}$. The branching ratio is given by
$\frac{\mathrm{Br}(\ell^{i}\to\ell^{j}\gamma)}{\mathrm{Br}(\ell^{i}\to\ell^{j}\nu_{\ell^{i}}\overline{\nu}_{\ell^{j}})}=\frac{3\alpha}{64\pi G_{F}^{2}}\left|\sum_{\alpha}\frac{(h^{\alpha}_{j})^{\ast}h^{\alpha}_{i}f(r_{\alpha})}{m_{S}^{2}}\right|^{2},$ (53)
where $r_{\alpha}=m_{N^{\alpha}}^{2}/m_{S}^{2}$, and
$f(x)=\frac{1-6x+3x^{2}+2x^{3}-6x^{2}\log x}{6(-1+x)^{4}}.$ (54)
Three body decays $\ell^{i}\to\ell^{j}\ell^{k}\overline{\ell^{m}}$ also occurs at one-loop level. There are two kinds of Feynman diagrams. One is the box diagram, and the other is the photon penguin diagram. As discussed below, we are interested in parameter regions where the size of $h_{i}^{\alpha}$ is relatively large. In such a case, the box diagram gives the dominant contribution. Thus, we neglect the contribution from the photon penguin diagram in the following.
The branching ratio for $\ell^{i}\to\ell^{j}\ell^{k}\overline{\ell^{m}}$ is given by
$\frac{\mathrm{Br}(\ell^{i}\to\ell^{j}\ell^{k}\overline{\ell^{m}})}{\mathrm{Br}(\ell^{i}\to\ell^{j}\nu_{\ell^{i}}\overline{\nu}_{\ell^{j}})}=\frac{2-\delta_{jk}}{4096\pi^{4}G_{F}^{2}}\bigl{|}A_{i,m}^{jk}\bigr{|}^{2},$ (55)
where $A_{i,m}^{jk}=B_{i,m}^{jk}+B_{i,m}^{kj}+C_{i,m}^{jk}+C_{i,m}^{kj}$, and the tensors $B_{i,m}^{jk}$ and $C_{i,m}^{jk}$ are defined as
$B_{i,m}^{jk}=\frac{h^{\alpha}_{i}h^{\beta}_{m}(h^{\alpha}_{j}h^{\beta}_{k})^{\ast}}{m_{S}^{2}}\left(\frac{g_{2}\left(r_{\alpha}\right)-g_{2}\left(r_{\beta}\right)}{r_{\alpha}-r_{\beta}}\right),$ (56)
$C_{i,m}^{jk}=\frac{h^{\alpha}_{i}h^{\alpha}_{m}(h^{\beta}_{j}h^{\beta}_{k})^{\ast}}{m_{S}^{2}}\sqrt{r_{\alpha}r_{\beta}}\left(\frac{g_{1}(r_{\alpha})-g_{1}(r_{\beta})}{r_{\alpha}-r_{\beta}}\right),$ (57)
by using the function $g_{k}(x)$ ($k=1,2$) given by
$g_{k}(x)=\frac{1}{2^{k}}\Bigl{\{}\frac{x^{k}}{(1-x)^{2}}\log x+\frac{1}{1-x}\Bigr{\}}.$ (58)
III.3 dark matter
In the model, there are two dark matter candidates: the real scalar boson $\eta$ and the lightest of the Majorana fermion $N^{\alpha}$. Since $N^{\alpha}$ is favored to be heavier than $S^{\pm}$ to avoid the constraint from the lepton flavor violating decays as discussed in Sec. V, we consider the case that $\eta$ is the dark matter particle.
The relic abundance of dark matter is determined by the thermally averaged cross section $\sigma v\right>$, which is a function of the temperature $T$ evaluated by
$\sigma v\right>=\frac{2x}{K_{2}(x)^{2}}\int_{1}^{\infty}\mathrm{d}y\ \sigma_{\text{\sf all}}vy\sqrt{y-1}K_{1}(2x\sqrt{y}),$ (59)
where $x=m_{\eta}/T$, $y=s/(4m_{\eta}^{2})$ with the Mandelstam variable $s$, $\sigma_{\text{\sf all}}$ is the sum of all pair-annihilation processes of $\eta$, and $K_{1}(x)$ and $K_{2}(x)$ are the modified Bessel functions of the second kind. Here, we consider pair-annihilations of $\eta$ into a pair of the SM fermions ($f^{i}\bar{f^{i}}$), the weak boson’s pair ($W^{+}W^{-}$ and $ZZ$), diphoton ($\gamma\gamma$) and a pair of the additional Higgs bosons ($H_{i}H_{j}$ and $H^{+}H^{-}$). The former three have already been investigated in the previous works Aoki:2008av ; Aoki:2011zg ; Aoki:2022bkg , whereas the last one is a new update in this paper. We show the formulas for $\sigma v$ of these annihilation channels in the following.
The fermionic channel $\eta\eta\to f^{i}\bar{f^{i}}$, where $f=u$, $d$, $\ell$ and $i=1,2,3$, is mediated by $H_{a}^{(\ast)}$ and is generated by the Yukawa interaction and the following three-point scalar interactions;
$\mathcal{L}=-\sum_{a}\frac{v}{2}\lambda_{\eta\eta a}H_{a}\eta\eta,$ (60)
where
$\lambda_{\eta\eta a}=(\sigma_{1}c_{\beta}^{2}+\sigma_{2}s_{\beta}^{2})R_{1a}-(\sigma_{1}-\sigma_{2})s_{\beta}c_{\beta}R_{2a}.$ (61)
We note that there is no term proportional to $R_{3a}$ because a pair of $\eta$ does not couple to the $CP$-odd component $h_{3}^{\prime}$ at tree level. This is a different point from Ref. Aoki:2022bkg which is also a $CP$-violating extension of the original AKS model. The cross section is given by
$(\sigma v)_{f^{i}}=$ $\frac{N_{c}^{f}m_{f^{i}}^{2}}{4\pi}\Biggl{\{}h(m_{f^{i}})^{3}\left|\sum_{a=1}^{3}\frac{\lambda_{\eta\eta a}Z_{f}^{a}}{s-m_{H_{a}}^{2}+im_{H_{a}}\Gamma_{H_{a}}}\right|^{2}+h(m_{f^{i}})\left|\sum_{a=1}^{3}\frac{\lambda_{\eta\eta a}X_{f}^{a}}{s-m_{H_{a}}^{2}+im_{H_{a}}\Gamma_{H_{a}}}\right|^{2}\Biggr{\}},$ (62)
where $N_{c}^{f}$ is the color factor of the fermion, $h(x)=\sqrt{1-4x^{2}/s}$, $\Gamma_{H_{a}}$ is the decay rate of $H_{a}$, and $Z_{f}^{a}$ and $X_{f}^{a}$ is the coupling constant of the scalar and pseudos scalar interaction between $H_{a}$ and the SM fermions given by
$Z_{u}^{a}=R_{1a}+R_{2a}\cot\beta,\quad X_{u}^{a}=-R_{3a}\cot\beta,$ (63)
$Z_{d}^{a}=R_{1a}+R_{2a}\cot\beta,\quad X_{d}^{a}=R_{3a}\cot\beta,$ (64)
$Z_{\ell}^{a}=R_{1a}+R_{2a}\tan\beta,\quad X_{\ell}^{a}=R_{3a}\tan\beta.$ (65)
The annihilations into the weak bosons $\eta\eta\to W^{+}W^{-}/ZZ$ are also mediated by $H_{a}^{(\ast)}$ and are generated by the weak interaction and the above $\eta\eta H_{a}$ coupling. The cross section is given by
$(\sigma v)_{V}=$ $\frac{S_{V}(\sigma_{1}c_{\beta}^{2}+\sigma_{2}s_{\beta}^{2})^{2}}{16\pi s}h(m_{V})\Bigl{\{}(s-2m_{V}^{2})^{2}+8m_{V}^{4}\Bigr{\}}\biggl{|}\sum_{a=1}^{3}\frac{R_{1a}}{(s-m_{H_{a}}^{2})^{2}+m_{H_{a}}^{2}\Gamma_{H_{a}}^{2}}\biggr{|}^{2},$ (66)
where $V=W$ or $Z$, and $S_{V}$ is two (one) for $V=W$ ($Z$). We note that there is no term proportional to $R_{2a}$ or $R_{3a}$ because $\Phi_{2}$ does not obtain the VEV, and $h_{2}^{\prime}$ and $h_{3}^{\prime}$ do not have three-point interaction $h_{a}^{\prime}VV$.
The diphoton annihilation is generated by two kinds of one-loop Feynman diagrams Aoki:2008av ; Aoki:2022bkg . The first one includes circle and triangle diagrams constituted by a loop of $H^{\pm}$ or $S^{\pm}$ like the scalar contribution in the Higgs diphoton decay with the replacement of the external Higgs boson into two external lines $\eta\eta$. The second one includes box diagrams constituted by both $H^{\pm}$ and $S^{\pm}$. The cross section is given by
$(\sigma v)_{\gamma}=$ $\frac{\alpha^{2}}{256\pi^{3}s}\int_{-1}^{1}\mathrm{d}\cos\theta\Bigl{\{}s^{2}|A_{\gamma\gamma}|^{2}+2s^{2}m_{\eta}^{2}\mathrm{Re}[A_{\gamma\gamma}B_{\gamma\gamma}^{\ast}]+|B_{\gamma\gamma}|^{2}\Bigl{(}(m_{\eta}^{4}-ut)^{2}+s^{2}m_{\eta}^{4}\Bigr{)}\Bigr{\}},$ (67)
where $\theta$ is the scattering angle in the CM frame, and $t$ and $u$ are the Mandelstam variables. See Ref. Aoki:2022bkg for the definition of $A_{\gamma\gamma}$ and $B_{\gamma\gamma}$.
Finally, we consider the annihilations into a pair of the Higgs bosons. The dominant contribution comes from four-point interactions $\eta\eta H_{a}H_{b}$ and $\eta\eta H^{+}H^{-}$ unless $m_{\eta}$ is very close to a half of the mass of the additional Higgs bosons $H_{a}$, and the decay $H_{a}\to H_{b}H_{c}$ ($b,c\neq a$) or $H_{a}\to H^{+}H^{-}$ is kinematically allowed. Thus, we consider only the contribution from the four-point interactions. All such four-point interactions are generated by the coupling $\sigma_{1}$ and $\sigma_{2}$. It is convenient to use the interactions in the Higgs basis;
$\mathcal{L}=-\frac{\eta^{2}}{2}\Bigl{\{}\sigma_{1}^{\prime}|\Phi_{1}|^{2}+\sigma_{2}^{\prime}|\Phi_{2}|^{2}+\sigma_{12}^{\prime}\bigl{(}\Phi_{1}^{\dagger}\Phi_{2}+\mathrm{h.c.}\bigr{)}\Bigr{\}},$ (68)
where
$\sigma_{1}^{\prime}=\sigma_{1}c_{\beta}^{2}+\sigma_{2}s_{\beta}^{2},\quad\sigma_{2}^{\prime}=\sigma_{1}s_{\beta}^{2}+\sigma_{2}c_{\beta}^{2},\quad\sigma_{12}^{\prime}=(\sigma_{2}-\sigma_{1})s_{\beta}c_{\beta}.$ (69)
Then, the four-point interaction is given by
[…]
$+\sigma_{2}^{\prime}\Bigl{(}2|H^{+}|^{2}+(R_{2a}R_{2b}+R_{3a}R_{3b})H_{a}H_{b}\Bigr{)}\Bigr{\}}.$ (70)
In particular, in the case of the small mixing ($R_{ab}\simeq\delta_{ab}$), it is approximated by
$\mathcal{L}\simeq-\frac{\eta^{2}}{4}\Bigl{\{}\sigma_{1}^{\prime}R_{1a}R_{1b}H_{1}^{2}+2\sigma_{12}^{\prime}H_{1}H_{2}+\sigma_{2}^{\prime}\Bigl{(}2|H^{+}|^{2}+H_{2}^{2}+H_{3}^{2}\Bigr{)}\Bigr{\}}.$ (71)
The cross section is given by
$(\sigma v)_{H}$ $=\frac{\sigma_{2}^{\prime}}{8\pi s}h(m_{H^{\pm}}^{2})+\sum_{a\geq b}\frac{\lambda_{\eta\eta ab}^{2}s_{ab}}{8\pi s}\tilde{h}(m_{H_{a}},m_{H_{b}}),$ (72)
where $\tilde{h}(x,y)=\sqrt{1-2(x^{2}+y^{2})/s+(x^{2}-y^{2})^{2}/s^{2}}$, $s_{ab}=1/(1+\delta_{ab})$, and $\lambda_{\eta\eta ab}$ is the vertex factor of $\eta\eta H_{a}H_{b}$ given by
$\lambda_{\eta\eta ab}=-\frac{1}{2s_{ab}}\Bigl{(}$ $\sigma_{1}^{\prime}R_{1a}R_{1b}+2\sigma_{12}^{\prime}R_{1a}R_{2b}+\sigma_{2}^{\prime}(R_{2a}R_{2b}+R_{3a}R_{3b})\Bigr{)}.$ (73)
IV Electric dipole moments
In this section, we investigate the eEDM $d_{e}$ in the model. There are two kinds of $CP$-violating phases in the model, those in the Yukawa matrix $h_{i}^{\alpha}$ and $\theta_{5}$ in the Higgs potential. They independently induce the eEDM, and they are denoted by $d_{e}^{h}$ and $d_{e}^{\theta_{5}}$ in the following, respectively.
First, we consider $d_{e}^{h}$. The leading contribution comes from two-loop diagrams of the quartic orders of the Yukawa matrix $h$, which are shown in Fig. 2. Diagrams in Fig. 2(a) give finite contributions to the EDM; on the other hand, contributions from diagrams in Fig. 2(b) are canceled although each of them is nonzero. The same cancellation was discussed in the scotogenic model in Ref. Abada:2018zra , and Ref. Fujiwara:2021vam has proved the cancellation mechanism with a more general setup by using the Ward-Takahashi identity.
(a) Feynman diagrams giving finite contributions
(b) Feynman diagrams with canceling contributions
Figure 2: Feynman diagrams for the EDM of the charged leptons via the Yukawa matrix $h$.
As a result, $d_{e}^{h}$ at this order is given by
$d_{e}^{h}=\frac{em_{e}}{m_{S}^{2}}\left(\frac{1}{16\pi^{2}}\right)^{2}\sum_{\alpha,\beta,j}I_{ij}^{\alpha\beta}\sqrt{r_{\alpha}r_{\beta}}\Bigl{(}\mathcal{J}^{1}_{\alpha\beta}+\mathcal{J}^{2}_{\alpha\beta}\Bigr{)},$ (74)
where $I_{ij}^{\alpha\beta}=\mathrm{Im}[(h_{i}^{\beta}h_{j}^{\beta})^{\ast}(h_{i}^{\alpha}h_{j}^{\alpha})]$, and the function $J_{\alpha\beta}^{1}$ and $J_{\alpha\beta}^{2}$ are given by the following integrals
$\mathcal{J}^{1}_{\alpha\beta}=\int\frac{stz\bigl{(}-(1-y)(1-z)+sy(1-y)+tyz\bigr{)}}{\bigl{[}y(1-y)sr_{\beta}+txr_{\alpha}+tz+uy(1-y)\bigr{]}^{2}},$ (75)
$\mathcal{J}^{2}_{\alpha\beta}=\int\frac{sty(1-y)\bigl{(}z-(1-y)s-zt\bigr{)}}{\bigl{[}y(1-y)sr_{\beta}+txr_{\alpha}+tz+uy(1-y)\bigr{]}^{2}},$ (76)
with
$\int$=∫ = $\int_{0}^{1}\mathrm{d}x\mathrm{d}y\mathrm{d}z\,\delta(1-x-y-z)\int_{0}^{1}\mathrm{d}s\mathrm{d}t\mathrm{d}u\,\delta(1-s-t-u).$ (77)
The loop functions $\mathcal{J}_{\alpha\beta}^{1}$ and $\mathcal{J}_{\alpha\beta}^{2}$ correspond to the sum of the upper two diagrams and the lower diagram in Fig. 2(a), respectively.
Here, we show some comments about comparisons with Ref. Abada:2018zra studying the eEDM in the scotogenic model Ma:2006km , which is described by Feynman diagrams with the same topology. We have numerically checked $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ are the same with the loop functions $I_{1+2}^{M}$ and $I_{3}^{M}$ defined in Ref. Abada:2018zra , respectively although their integral representations are different. The overall signs are opposite comparing Eq. (74) with their result. This is due to the different chirality of the charged leptons. In the scotogenic model, the left-handed leptons have the Yukawa interaction, while the right-handed ones do in our model. It induces the opposite signs of the coefficients of $\gamma_{5}$.
Next, we discuss the contribution from the Higgs sector $d_{e}^{\theta_{5}}$. This is the same with the eEDM in the THDMs studied in Refs. Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb . It is known that one-loop contribution is tiny because of the mass suppression by $m_{e}$, and the leading contribution comes from Barr-Zee type diagrams at two-loop level Barr:1990vd . We employ the formula in Ref. Altmannshofer:2020shb , which includes all the two-loop diagrams and gives the gauge-invariant result.
V Numerical evaluations and benchmark scenarios
In this section, we show some benchmark scenarios and their numerical evaluations. Before a detailed discussion, we roughly describe our strategy to determine the benchmark scenarios. We first consider the constraint from the neutrino oscillation data. The Casas-Ibarra parametrization in Eq. (50) tells us the following preferred mass spectrum;
$m_{H^{\pm}},\ m_{S},\ m_{\eta}=\mathcal{O}(100)~{}\mathrm{GeV},\quad m_{N^{\alpha}}=\mathcal{O}(1)~{}\mathrm{TeV}.$ (78)
This hierarchy implies that the real scalar dark matter $\eta$ is strongly favored in our model.
Next, we determine the structure of the Yukawa matrix $h$ with a fixed mass spectrum to avoid the current constraint on the LFV decays. We choose the values of the orthogonal matrix $O$ and unobserved parameters in the active neutrino mass matrix: the lightest neutrino mass $m_{\mathrm{lightest}}$ and Majorana $CP$-violating phases $\alpha_{1}$ and $\alpha_{2}$.
$\eta$ is responsible for all the observed relic density of DM, which is made by the freeze-out mechanism. There are two ways to make the annihilation rate large enough to reproduce the observed value $\Omega_{\mathrm{DM}}h^{2}\simeq 0.12$ under the current severe constraint on the direct detection of DM. One is the resonance effect of the on-shell Higgs boson, which is called the Higgs portal DM scenario. The other one uses the annihilation into the additional Higgs bosons by contact scalar interactions, which is denoted by a heavy dark matter scenario in the following. These make two benchmark scenarios with different values of $m_{\eta}$.
The mass of the neutral additional Higgs bosons $H_{2}$ and $H_{3}$ is almost ineffective in the above discussions. We determine them so as to satisfy the theoretical requirements on the Higgs potential and to avoid the constraint on the $S$, $T$ and $U$ parameters. Other parameters in the Higgs potential are fixed to satisfy the theoretical constraints.
Finally, we fix the $CP$-violating phase $\theta_{5}$ to avoid the current constraint by the destructive interference with the $CP$-violating phases in $h_{i}^{\alpha}$, which is already determined to avoid the constraint on the LFV decays. We will show that $\theta_{5}$ can be $\mathcal{O}(1)$, which is necessary to generate sufficient baryon asymmetry by electroweak baryogenesis.
V.1 Neutrino mass and the size Yukawa couplings $h^{\alpha}_{i}$
As long as $h_{i}^{\alpha}$ employs the Casas-Ibarra parametrization in Eq. (50), it can explain the neutrino oscillation data. $h_{i}^{\alpha}$ is then given by
$h_{i}^{\alpha}=\frac{(16\pi^{2})^{3/2}}{|\kappa\tan\beta|}\sum_{a=1}^{3}\sqrt{|\Lambda_{\alpha}|}O_{\alpha a}\sqrt{m_{\nu^{a}}}U^{\ast}_{ia}m_{\ell^{i}}^{-1}.$ (79)
Thus, the size of $h_{i}^{\alpha}$ can be roughly estimated by
$|h_{i}^{\alpha}|\sim 0.1\times\frac{(16\pi^{2})^{3/2}}{|\kappa\tan\beta|}\frac{\sqrt{|\Lambda_{\alpha}|\bar{m}_{\nu}}}{m_{\ell^{i}}},$ (80)
where $\bar{m}_{\nu}$ are the typical size of the neutrino mass. We have assumed that the factor $O_{a\alpha}U^{\ast}_{ia}$ and the summation about $a$ in Eq. (79) all together make a coefficient of $\mathcal{O}(0.1)$ for simplicity. [^1] This correctness of this assumption depends on the $CP$-violating phases $\psi_{i}$ in $O_{a\alpha}$. The large $CP$-violating phases ($|\psi_{i}|\simeq\pi/2$) make the bound stronger. As far as $|\psi_{i}|$ is not too large, we can use this assumption. We note $|h_{i}^{\alpha}|$ is proportional to $m_{\ell^{i}}^{-1}$, and $N$’s strongly couple to electrons.
To keep the perturbativity, the size of $h_{i}^{\alpha}$ should not be too large. The large couplings also lead to difficulty in avoiding the constraint on the LFV processes. We thus demand that $|h_{i}^{\alpha}|$ be less than three. Since $h_{i}^{\alpha}\propto\sqrt{\Lambda_{\alpha}}$, this requirement constraints $\Lambda_{\alpha}$, i.e. the mass spectrum of the new particles. The bound on $\Lambda_{\alpha}$ is roughly estimated by
$|\Lambda_{\alpha}|\lesssim\frac{m_{\ell^{i}}^{2}|\kappa\tan\beta|^{2}}{(16\pi^{2})^{3}\bar{m_{\nu}}}.$ (81)
Eq. (81) provides the most significant constraint in the case of $m_{\ell^{i}}=m_{e}$, which is given by
$|\Lambda_{\alpha}|\lesssim 1~{}\mathrm{TeV}\times\biggl{(}\frac{|\kappa\tan\beta|}{40}\biggr{)}^{2}\biggl{(}\frac{\bar{m}_{\nu}}{0.01~{}\mathrm{eV}}\biggr{)}^{-1}.$ (82)
In Fig. 3, we show $\Lambda_{\alpha}$ as a function of $m_{S}$ and $m_{N^{\alpha}}$. The red, blue and green curves are contours of $\Lambda_{\alpha}=1.0~{}\mathrm{TeV}$, $0.5~{}\mathrm{TeV}$ and $0.25~{}\mathrm{TeV}$, respectively. $m_{H^{\pm}}$ and $m_{\eta}$ are set to be $m_{H^{\pm}}=390~{}\mathrm{GeV}$ and $m_{\eta}=63~{}\mathrm{GeV}$ ($250~{}\mathrm{GeV}$) on the solid (dashed) curves. The gray line shows $m_{S}=m_{N^{\alpha}}$.
In the case of $\kappa\tan\beta=40$ and $\bar{m}_{\nu}=0.01~{}\mathrm{eV}$, the upper bound on $|\Lambda_{\alpha}|$ is given by the red curves, and the bound is stronger for smaller $\kappa\tan\beta$. We can see that heavier $N^{\alpha}$ than $S^{\pm}$ are strongly preferred. This is because $\Lambda_{\alpha}$ is proportional to $m_{S}^{2}$ if $m_{S}\gg m_{N^{\alpha}}$, $m_{H^{\pm}}$, $m_{\eta}$, while it linearly depends on $m_{N^{\alpha}}$ if $N^{\alpha}$ is the heaviest particle. The hierarchy $m_{N^{\alpha}}>m_{S}$ implies that $\eta$ is the dark matter particle in the model.
Figure 3: $\Lambda_{\alpha}$ as a function of $m_{S}$ and $m_{N^{\alpha}}$. The red, blue and green curves correspond to $\Lambda_{\alpha}=1.0~{}\mathrm{TeV}$, $0.5~{}\mathrm{TeV}$ and $0.25~{}\mathrm{TeV}$, respectively. The masses of the other scalar bosons are set as $m_{H^{\pm}}=390~{}\mathrm{GeV}$ and $m_{\eta}=63~{}\mathrm{GeV}$ ($250~{}\mathrm{GeV}$) on the solid (dashed) curves. The gray line shows $m_{S}=m_{N^{\alpha}}$.
V.2 The LFV decays
Next, we consider the constraint from the LFV decays by using the Casas-Ibarra parametrization of $h_{i}^{\alpha}$. Ref. Toma:2013zsa performed a similar analysis in the scotogenic model Ma:2006km . For simplicity, we use an approximation such that all of $N$’s have the same mass $\bar{m}_{N}$. Even in the case that they have different masses, the result expects to give a good approximation by replacing $\bar{m}_{N}$ into their typical mass scale unless the mass difference is too large.
Since $h_{i}^{\alpha}\propto m_{\ell^{i}}^{-1}$, the LFV decays including the light leptons $e$ and $\mu$ expect to give significant constraints. Thus, we show only the analysis for $\mu\to e\gamma$ and $\mu\to 3e$. It is straightforward to apply the same method to the other LFV decays. By using Eqs. (53), (55) and (79), their branching ratios are evaluated as
$\mathrm{Br}(\mu\to e\gamma)=$ $\frac{3\alpha}{64\pi G_{F}^{2}m_{S}^{4}}\biggl{(}\frac{(16\pi^{2})^{3}\bar{\Lambda}F}{m_{\mu}m_{e}\kappa^{2}\tan^{2}\beta}\biggr{)}^{2}\bigl{|}X_{\mu e}\bigr{|}^{2}$
$\simeq$ $\,\bigl{(}4\times 10^{27}\bigr{)}\biggl{(}\frac{1}{\kappa\tan\beta}\biggr{)}^{4}\biggl{(}\frac{\bar{\Lambda}^{2}F^{2}}{m_{S}^{4}}\biggr{)}\bigl{|}X_{\mu e}\bigr{|}^{2},$ (83)
$\mathrm{Br}(\mu\to 3e)=$ $\frac{1}{1024\pi^{4}G_{F}^{2}m_{S}^{4}}\left(\frac{(16\pi^{2})^{6}\bar{\Lambda}^{2}G}{m_{\mu}m_{e}^{3}\kappa^{4}\tan^{4}\beta}\right)^{2}\bigl{|}X_{\mu e}X_{ee}-a(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}\bigr{|}^{2}$
$\simeq$ $\bigl{(}10^{5}~{}\mathrm{meV}^{-4}\bigr{)}\biggl{(}\frac{1}{\kappa\tan\beta}\biggr{)}^{8}\biggl{(}\frac{\bar{\Lambda}^{4}G^{2}}{m_{S}^{4}}\biggr{)}\bigl{|}X_{\mu e}X_{ee}-a(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}\bigr{|}^{2}.$ (84)
where $\bar{\Lambda}$ is $\Lambda_{\alpha}$ with $m_{N^{\alpha}}=\bar{m}_{N}$, $F$, $G$ and $a$ is defined as
$F\equiv f(\bar{r})=\frac{1-6\bar{r}+3\bar{r}^{2}+2\bar{r}^{3}-6\bar{r}^{2}\log\bar{r}}{6(1-\bar{r})^{4}},$ (85)
$G\equiv\frac{\mathrm{d}g_{2}(\bar{r})}{\mathrm{d}\bar{r}}=\frac{1}{4(1-\bar{r})^{2}}\Bigl{(}1+\bar{r}+\frac{2\bar{r}}{1-\bar{r}}\log\bar{r}\Bigr{)},$ (86)
$a\equiv-\frac{1}{\bar{r}}\frac{\mathrm{d}g_{2}(\bar{r})}{\mathrm{d}\bar{r}}\biggl{(}\frac{\mathrm{d}g_{1}(\bar{r})}{\mathrm{d}\bar{r}}\biggr{)}^{-1}=\frac{2\bar{r}(2\bar{r}-2-(1+\bar{r})\log\bar{r})}{1-\bar{r}^{2}+2\bar{r}\log\bar{r}},$ (87)
with $\bar{r}=\bar{m}_{N}^{2}/m_{S}^{2}$, and $X_{\ell^{i}\ell^{j}}$ is given by
$X_{\ell^{i}\ell^{j}}=\sum_{a,b,\alpha}\sqrt{m_{\nu^{a}}m_{\nu^{b}}}U_{ia}^{\ast}U_{jb}O_{\alpha a}O_{\alpha b}^{\ast}.$ (88)
$X_{\ell^{i}\ell^{j}}$ satisfies $X_{\ell^{i}\ell^{j}}=X_{\ell^{j}\ell^{i}}^{\ast}$ and has the mass dimension one. If the matrix $O$ is a real matrix, i.e., $\psi_{i}=0$ ($i=1,2,3$), $X_{\ell^{i}\ell^{j}}$ is independent of $O_{a\alpha}$ because $\sum_{\alpha}O_{\alpha a}O_{\alpha b}^{\ast}=\sum_{\alpha}O_{\alpha a}O_{\alpha b}=\delta_{ab}$. It is simplified as
$X_{\ell^{i}\ell^{j}}=(U^{\ast}DU^{\mathrm{T}})_{ij}=\sum_{a}m_{\nu^{a}}U_{ia}^{\ast}U_{ja}\quad\text{if all $O_{a\alpha}$ are real.}$. (89)
Figure 4: The upper bound on $|X_{\mu e}|$ and $\sqrt{|Y_{\mu e}|}$ as functions of $\bar{m}_{N}$ and $m_{S}$ in the case of $\kappa\tan\beta=30$. The red (blue) curves are contours of $|X_{\mu e}|^{\mathrm{up}}$ ($\sqrt{|Y_{\mu e}|^{\mathrm{up}}}$). The solid and dashed curves correspond to the cases of ($m_{H^{\pm}}$, $m_{\eta}$) = (390 GeV, 63 GeV) and (390 GeV, 250 GeV), respectively. The gray regions are unfavored by the bound in Eq. (82) in the former case. The dashed gray curve is the bound in the latter case.
By using the current upper limits, $\mathrm{Br}(\mu\to e\gamma)<3.1\times 10^{-13}$ MEGII:2023ltw and $\mathrm{Br}(\mu\to 3e)<1.0\times 10^{-12}$ SINDRUM:1987nra at $90~{}\%$ confidence level, we have
$\bigl{|}X_{\mu e}\bigr{|}<\bigl{(}0.01~{}\mathrm{meV}\bigr{)}\biggl{(}\frac{\kappa\tan\beta}{40}\biggr{)}^{2}\biggl{|}\frac{m_{S}^{2}/(\bar{\Lambda}F)}{1~{}\mathrm{TeV}}\biggr{|},$ (90)
$\bigl{|}Y_{\mu e}\bigr{|}\equiv\bigl{|}X_{\mu e}X_{ee}-a(M_{\nu})_{\mu e}(M_{\nu})^{\ast}_{ee}\bigr{|}<(0.1~{}\mathrm{meV})^{2}\biggl{(}\frac{\kappa\tan\beta}{40}\biggr{)}^{4}\biggl{(}\frac{m_{S}^{2}}{\bar{\Lambda}^{2}G}\biggr{)}.$ (91)
The upper bound on $|X_{\mu e}|$ and $|Y_{\mu e}|$, which are denoted as $|X_{\mu e}|^{\mathrm{up}}$ and $|Y_{\mu e}|^{\mathrm{up}}$ in the following, are functions of $\kappa\tan\beta$ and the masses of the new particles.
Fig. 4 shows $|X_{\mu e}|^{\mathrm{up}}$ and $\sqrt{|Y_{\mu e}|^{\mathrm{up}}}$ in the cases of $\kappa\tan\beta=30$. The red curves are contours for $|X_{\mu e}|^{\mathrm{up}}=1.0~{}\mathrm{meV}$ and $0.1~{}\mathrm{meV}$, and the blue ones are those for $\sqrt{|Y_{\mu e}|^{\mathrm{up}}}=1.0~{}\mathrm{meV}$ and $0.3~{}\mathrm{meV}$. The solid and dashed lines correspond to the cases of ($m_{H^{\pm}}$, $m_{\eta}$) = (390 GeV, 63 GeV) and (390 GeV, 250 GeV), respectively. The gray regions represent unfavored regions from the bound in Eq. (82) for ($m_{H^{\pm}}$, $m_{\eta}$) = (390 GeV, 63 GeV), and the gray dashed curve is the bound for ($m_{H^{\pm}}$, $m_{\eta}$) = (390 GeV, 250 GeV). The bounds are stronger for smaller $\kappa\tan\beta$ because the Yukawa couplings $h^{\alpha}_{i}$ are proportional to $|\kappa\tan\beta|^{-1}$.
Table 2: Input values for neutrino mass matrix PDG
$\sin^{2}\theta_{12}$ $\sin^{2}\theta_{23}$ $\sin^{2}\theta_{13}$ $m_{\nu^{2}}^{2}-m_{\nu^{1}}^{2}$ ($\mathrm{eV^{2}}$) $m_{\nu^{3}}^{2}-m_{\nu^{2}}^{2}$ ($\mathrm{eV^{2}}$)
NO $0.307$ $0.534$ $2.20\times 10^{-2}$ $7.53\times 10^{-5}$ $2.437\times 10^{-3}$
IO $0.547$ $-2.519\times 10^{-3}$
(a) Normal Ordering
(b) Inverted Ordering
Figure 5: Possible values of $|(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}|~{}\mathrm{meV^{2}}$ as a function of $m_{\mathrm{lightest}}$. The darker-colored lines represent values without $CP$-violating phases ($\delta=\alpha_{1}=\alpha_{2}=0$). The lighter-colored regions are possible values with $CP$-violating phases in the regions $\delta\in[1.02\pi,1.44\pi]$, $\alpha_{1,2}\in[0,2\pi)$.
Thus, we roughly evaluate the constraint as $|X_{\mu e}|<1.0~{}\mathrm{meV}$ and $\sqrt{|Y_{\mu e}|}<1.0~{}\mathrm{meV}$. For the latter equation, we consider the sufficient condition $\sqrt{|X_{\mu e}X_{ee}|}<1.0~{}\mathrm{meV}$ and $\sqrt{|a(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}|}<1.0~{}\mathrm{meV}$. In addition, in the case of $10m_{S}<\bar{m}_{N}<100m_{S}$, $a$ is given by $165<a<$ 16. As a result, we employ the following rough conditions to avoid the LFV constraints [^2] Even if $|(M_{\nu})_{\mu e}(M_{\nu})^{\ast}_{ee}|$ and $|X_{\mu e}X_{ee}|$ do not satisfy these bounds, it would be possible to make $|Y_{\mu e}|$ small enough to avoid the constraint from the LFV decays by a large cancellation between them. Here, we do not consider such a case.
$\sqrt{|(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}|}<0.25~{}\mathrm{meV},\quad|X_{\mu e}|<1.0~{}\mathrm{meV},\quad\sqrt{|X_{\mu e}X_{ee}|}<1.0~{}\mathrm{meV}.$ (92)
(a) Normal Ordering
(b) Inverted Ordering
Figure 6: $|X_{\mu e}|~{}\mathrm{meV}$ for the normal ordering and inverted ordering neutrino mass data. On the darkest color points, $\psi_{i}$ ($i=1,2,3$) are zero. The middle dark and light color points mean that $\psi_{i}$ are in the range $[0,0.02]$ and $[0.02,0.2]$, respectively. The solid black line is the upper limit on $m_{\mathrm{lightest}}$ from the CMB data. The dashed black line represents $|X_{\mu e}|=1~{}\mathrm{meV}$.
In Fig. 5, we show $\sqrt{|(M_{\nu})_{\mu e}(M_{\nu})_{ee}^{\ast}|}$ as a function of the lightest neutrino mass $m_{\mathrm{lightest}}$ for (a) normal ordering neutrino masses (NO) and (b) Inverted ordering neutrino masses (IO). We used the data from Particle Data Group PDG in Table 2 to make the figure. We neglect their small uncertainties. The dashed black lines represent the upper limit in Eq. (92), and the solid black lines correspond to the cosmological bound $\sum_{i=1}^{3}m_{\nu^{i}}<0.12~{}\mathrm{eV}$ Planck:2018vyg in each case, where heavier $m_{\mathrm{lightest}}$ is excluded. The darker-colored lines represent the value without $CP$-violating phases, i.e. $\delta=\alpha_{1}=\alpha_{2}=0$. The lighter-colored regions represent the possible values with switching on the $CP$-violating phases in the regions $\delta\in[1.02\pi,1.44\pi]$, $\alpha_{1,2}\in[0,2\pi)$, where $\delta$ is in the $1~{}\sigma$ region of the global fit data in PDG PDG . We can see that non-zero $CP$-violating phases are necessary to satisfy the condition in Eq. (92). Their interferences can make the value much smaller than the prediction in the $CP$-conserving case. In particular, in the inverted ordering case, we need a stronger interference effect to avoid the constraint.
(a) Normal Ordering
(b) Inverted Ordering
Figure 7: $|X_{\mu e}X_{ee}|~{}\mathrm{meV}$ for the normal ordering and inverted ordering neutrino mass data. The meaning of colors and black lines is the same as Fig. 6.
Next, we show possible values of $|X_{\mu e}|$ and $\sqrt{|X_{\mu e}X_{ee}|}$. These depend on the orthogonal matrix $O_{\alpha a}$ and have more parameter than $\sqrt{|(M_{\nu})_{e\mu}(M_{\nu})_{ee}^{\ast}|}$. Thus, we show scatter plots for each quantity, not a whole region of possible values, to reduce the cost of the calculation. We assume that the rotation angles $\gamma_{i}$ ($i=1$, 2, 3) are in the region $[0,2\pi)$. For $CP$-violating phases $\psi_{i}$ ($i=1$, 2, 3), we consider three cases (i) $\psi_{i}=0$, (ii) $|\psi_{i}|\in[0,0.02]$ and (iii) $|\psi_{i}|\in[0.02,0.2]$. Although $\psi_{i}$ can take a value in the region $[-\pi/2,\pi/2]$ in principle, we consider only these small $|\psi_{i}|$ regions because large $|\psi_{i}|$ make the Yukawa couplings $|h_{i}^{\alpha}|$ huge. The results for $|X_{\mu e}|$ and $\sqrt{|X_{\mu e}X_{ee}|}$ are shown in Figs. 6 and 7, respectively. The green (red) points are data for the normal (inverted) ordering neutrino masses. The darker-, middle-, and lighter-colored points correspond to the cases (i), (ii) and (iii), respectively.
We can see that in all the figures, the condition in Eq. (92) cannot be satisfied only with the $CP$-violating phases in the PMNS matrix, and non-zero $\psi_{i}$ are necessary. In case (ii), distributions of the points are broader; however, there are few points avoiding constraints. Consequently, the condition in Eq. (92) requires $|\psi_{i}|\gtrsim 0.02$. In case (iii) many allowed points appear in the figure for $|X_{\mu e}|$. Although the number of allowed points is smaller in the figure for $\sqrt{|X_{\mu e}X_{ee}|}$, we can also find some allowed points.
Before closing the section about the constraint on the LFV decays, we comment on the uncertainties of the above analysis. As mentioned at the beginning of this section, we assumed the degenerated $m_{N^{\alpha}}$. If the mass differences $\Delta m_{N}$ are switched on, there would be deviations with the order of $\Delta m_{N}/\bar{m}_{N}$. In addition, we roughly estimated the factors of the equations. Thus, we expect the $\mathcal{O}(100)~{}\%$ uncertainties for each value, e.g., the upper values in Eq. (92). Despite this fact, it is an important observation that non-zero $CP$-violating phases are necessary not only in the PMNS matrix but also in the matrix $O_{\alpha a}$ to avoid the constraint on the LFV decays. This conclusion would not change even in a more detailed analysis because the rough criteria in Eq. (92) and the $CP$-conserving values are different by many orders of magnitude. Points in Figs. 6 and 7 would be a good first candidate in searching for benchmark scenarios.
V.3 Dark matter
As explained above, $m_{N^{\alpha}}>m_{S}$ is preferred to explain the neutrino oscillation data and avoid the constraint on the LFV processes. Thus, $\eta$ is the dark matter particle in this model. The freeze-out mechanism can produce the observed relic density. Such a dark matter particle is strongly constrained by direct searches of the dark matter nucleus scattering XENON:2018voc ; PandaX-4T:2021bab ; LZ:2022lsv . We employ the latest result of the LZ collaboration LZ:2022lsv as a constraint.
The spin-independent cross section for $\eta N\to\eta N$, where $N$ is a nucleon, is approximately given by
$\sigma_{\mathrm{SI}}\simeq\frac{g^{2}m_{N}^{2}v^{2}}{4\pi(m_{\eta}+m_{N})^{2}m_{H_{1}}^{4}}(\sigma_{1}\cos\beta+\sigma_{2}\sin\beta)^{2},$ (93)
where $m_{N}\simeq 1~{}\mathrm{GeV}$ is the nucleon mass, $g\simeq 10^{-3}$ Shifman:1978zn is the coupling constant of the nucleon-Higgs interaction. In Eq. (93), we neglect the tiny mixing among the neutral Higgs bosons and the interaction between the nucleons and $H_{2,3}$, which is suppressed by $\cot\beta$. Since $\sigma_{1}\cos\beta+\sigma_{2}\sin\beta\simeq\sigma_{2}$ for large $\tan\beta$, $\sigma_{\mathrm{SI}}$ is almost independent of $\sigma_{1}$ and is a function of $m_{\eta}$ and $\sigma_{2}$. Fig. 8 shows the excluded region by the latest LZ result on the $m_{\eta}$$|\sigma_{2}|$ plane.
Figure 8: The constraint on $m_{\eta}$$|\sigma_{2}|$ plane from the latest constraint on the direct detection.
Under this constraint, there are two scenarios to produce the observed relic density: (i) Higgs portal DM scenario and (ii) Heavy DM scenario. In the Higgs portal DM scenario, we assume $m_{\eta}\simeq m_{H_{1}}/2\simeq 63~{}\mathrm{GeV}$. Although the direct-detection constraint is severe for such a dark matter, the resonance effect of the on-shell Higgs boson can enhance the annihilation cross section $\eta\eta\to H_{1}\to f\bar{f}(gg)$, and the correct relic density can be reproduced with $|\sigma_{2}|$ tiny enough to avoid the constraint.
In the heavy DM scenario, we assume relatively heavy $\eta$ whose pair can be annihilated into a pair of the additional Higgs bosons $H_{i}H_{j}$ and/or $H^{+}H^{-}$. Neglecting small mixing among neutral Higgs bosons and large $\tan\beta$, the contact interactions between $\eta$ and the additional Higgs bosons are given by
$\mathcal{L}_{\eta\eta H}\simeq-\frac{\sigma_{1}}{2}\Bigl{\{}|H^{\pm}|^{2}+\frac{1}{2}\bigl{(}H_{2}^{2}+H_{3}^{2}\bigr{)}\Bigr{\}}.$ (94)
[…]
V.4 Benchmark scenarios
According to the above analysis, we propose the following two benchmark scenarios, Scenario I and Scenario II. In both scenarios, the neutrinos have the normal ordering mass hierarchy. It would be possible to find a scenario for the inverted mass hierarchy by using the above discussions although we do not show it in this paper.
V.4.1 Scenario I
In Scenario I, we assume the following input values for $M_{h}^{2}$ in both benchmark scenarios.
$M=270~{}\mathrm{GeV},\quad\tilde{m}_{A}=360~{}\mathrm{GeV},\quad\tilde{m}_{H}^{2}-M^{2}=300~{}\mathrm{GeV^{2}},$
$\tan\beta=18,\quad s_{\beta-\tilde{\alpha}}=1.0,\quad\theta_{5}=-0.315\pi\simeq-0.990.$ (95)
$\tilde{m}_{h}$ is determined to obtain the observed Higgs boson mass $H_{1}=125~{}\mathrm{GeV}$ and is given by $\tilde{m}_{h}\simeq 126~{}\mathrm{GeV}$.
These input values lead to the following mass eigenvalues and the mixing matrix;
$m_{H_{1}}\simeq 125~{}\mathrm{GeV},\quad m_{H_{2}}\simeq 224~{}\mathrm{GeV},\quad m_{H_{3}}\simeq 391~{}\mathrm{GeV},$
$R\simeq\begin{pmatrix}0.997&0.0657&-0.0307\\ -0.0433&0.880&0.474\\ 0.0581&-0.471&0.880\\ \end{pmatrix}.$
Since $|R_{11}-1|\simeq 0.003$, it is expected to cause $\mathcal{O}(0.1)~{}\%$ deviations in the Higgs couplings from the SM prediction, which is small enough to satisfy the current LHC data.
$m_{H^{\pm}}$ is degenerated to $m_{H_{3}}$ to avoid the constraint from oblique paramters Peskin:1990zt , in particular the $T$ parameter;
$m_{H^{\pm}}=m_{H_{3}}\simeq 391~{}\mathrm{GeV}.$ (96)
In the limit $m_{H^{\pm}}=m_{H_{3}}$, the custodial symmetry of the Higgs potential Sikivie:1980hm ; Haber:1992py ; Pomarol:1993mu ; Gerard:2007kn ; Haber:2010bw ; Grzadkowski:2010dj ; Aiko:2020atr is restored and the new physics contribution vanishes. We have also evaluated the $S$ and $U$ parameters by using formulas in Ref. Branco:2011iw . The result is given by
$S\simeq-0.011,\quad T\simeq 0,\quad U\simeq-1.9\times 10^{-5}.$ (97)
The current result of the electroweak global fit gives $S=-0.01\pm 0.07$ and $T=0.04\pm 0.06$ with $U=0$ PDG . Thus, the benchmark scenario satisfies this constraint. We note that the other new particles do not contribute to the $STU$ parameters because they are the isospin singlets.
The original scalar quartic couplings in $V_{\mathrm{THDM}}$ is evaluated by using the above input values as follows;
$\begin{array}[]{l}\lambda_{1}\simeq 1.87,\quad\lambda_{2}\simeq 0.263,\quad\lambda_{3}\simeq 2.91,\\[5.0pt] \lambda_{4}\simeq-1.71,\quad\lambda_{5}^{R}\simeq-0.937,\quad\lambda_{5}^{I}\simeq 1.43.\\ \end{array}$ (100)
We have checked that these quartic couplings satisfy the conditions for the vacuum stability in Eqs. (27)–(28) and those for the perturbative unitarity in Eqs. (32)–(37). In addition, we expect that they are small enough to satisfy the triviality bound up to $10~{}\mathrm{TeV}$ Kanemura:2023wap .
Next, we consider the $Z_{2}$-odd particles. We assume the following mass spectrum;
$m_{S}=325~{}\mathrm{GeV},\quad m_{\eta}=63~{}\mathrm{GeV},$
$(m_{N^{1}},m_{N^{2}},m_{N^{3}})=(2500,3000,3500)~{}\mathrm{GeV}.$ (101)
We note that $m_{\eta}\simeq m_{H_{1}}/2$, and the DM relic density is explained in the Higgs portal scenario as discussed below.
The new physics scales in neutrino mass, $\Lambda_{\alpha}$ ($\alpha=1$, 2, 3), are determined by $m_{H^{\pm}}$, $m_{S}$, $m_{\eta}$ and $m_{N^{\alpha}}$ as follows [^4] We have used the Monte Carlo integration in evaluating loop integrals in $\Lambda_{\alpha}$ with $10^{9}$ sample points and have checked the uncertainties are smaller than $\mathcal{O}(0.1)~{}\%$.
$(\Lambda_{1},\Lambda_{2},\Lambda_{3})\simeq(269,271,276)~{}\mathrm{GeV}.$ (102)
In addition, we assume the following input values for the other parameters necessary for determining $h_{i}^{\alpha}$;
$\kappa\tan\beta=30,\quad m_{\nu^{1}}=4.76~{}\mathrm{meV},\quad\delta=1.12\pi,\quad\alpha_{1}=1.51\pi,\quad\alpha_{2}=1.12\pi,$
$\gamma_{1}=0.866\pi,\quad\gamma_{2}=0.861\pi,\quad\gamma_{3}=1.56\pi,\quad\psi_{1}=-0.0971,\quad\psi_{2}=0,\quad\psi_{2}=0.109\pi.$ (103)
Then, $h_{i}^{\alpha}$ are given by
$\begin{pmatrix}h^{1}_{e}&h^{1}_{\mu}&h^{1}_{\tau}\\ h^{2}_{e}&h^{2}_{\mu}&h^{2}_{\tau}\\ h^{3}_{e}&h^{3}_{\mu}&h^{3}_{\tau}\\ \end{pmatrix}\simeq\begin{pmatrix}2.85e^{0.955\pi i}&0.0103e^{0.386\pi i}&0.00242e^{-0.0717\pi i}\\ 2.31e^{-0.438\pi i}&0.0188e^{-0.928\pi i}&0.00208e^{-0.728\pi i}\\ 2.04e^{-0.480\pi i}&0.0543e^{-0.0291\pi i}&0.00221e^{-0.0828\pi i}\\ \end{pmatrix},$ (104)
where we used the neutrino oscillation data in the normal ordering case in Table 2. We can see the hierarchy $|h_{e}^{\alpha}|\gg|h_{\mu}^{\alpha}|>|h_{\tau}^{\alpha}|$ which is expected from the Casas-Ibarra parametrization, $h^{\alpha}_{i}\propto m_{\ell^{i}}^{-1}$.
Table 3: Branching ratios for LFV decays in Scenario I
Predictions Upper bounds
$\mu\to e\gamma$ $2.95\times 10^{-14}$ $3.1\times 10^{-13}$ MEGII:2023ltw
$\tau\to e\gamma$ $4.73\times 10^{-15}$ $3.3\times 10^{-8}$ BaBar:2009hkt
$\tau\to\mu\gamma$ $2.49\times 10^{-18}$ $4.4\times 10^{-8}$ BaBar:2009hkt
Predictions Upper bounds
$\mu\to 3e$ $4.68\times 10^{-13}$ $1.0\times 10^{-12}$ SINDRUM:1987nra
$\tau\to 3e$ $4.84\times 10^{-10}$ $2.7\times 10^{-8}$ Hayasaka:2010np
$\tau\to 3\mu$ $4.88\times 10^{-20}$ $2.1\times 10^{-8}$ Hayasaka:2010np
Predictions Upper bounds
$\tau\to e\mu\overline{e}$ $1.14\times 10^{-16}$ $1.8\times 10^{-8}$ Hayasaka:2010np
$\tau\to\mu\mu\overline{e}$ $5.77\times 10^{-17}$ $1.7\times 10^{-8}$ Hayasaka:2010np
$\tau\to ee\overline{\mu}$ $1.46\times 10^{-13}$ $1.5\times 10^{-8}$ Hayasaka:2010np
$\tau\to e\mu\overline{\mu}$ $1.14\times 10^{-16}$ $2.7\times 10^{-8}$ Hayasaka:2010np
This Yukawa interaction $h_{i}^{\alpha}$ induces the LFV decays as discussed in Sec. III.2. Their branching ratios and current upper bounds are summarized in Table 3. The current constraints can be avoided for all processes.
The DM relic density can be explained by the freeze-out mechanism through the annihilation $\eta\eta\to H_{1}\to f\bar{f}$. As explained above, the annihilation cross section can be large enough due to the resonance effect of $H_{1}$. $|\sigma_{2}|$ determines the relic density and is assumed as
$\sigma_{2}=1.1\times 10^{-3}.$ (105)
This and $m_{\eta}=63~{}\mathrm{GeV}$ reproduce the observed relic density $\Omega_{\mathrm{DM}}h^{2}=0.12$ Planck:2018vyg . In addition, we can see that the direct detection constraint can avoided in Fig. 8.
Finally, we evaluate the eEDM. By using Eqs. (104) and (V.4.1), the contribution from $h_{i}^{\alpha}$ is evaluated as
$d_{e}^{h}\simeq-5.82\times 10^{-29}~{}\mathrm{e\ cm}.$ (106)
On the other hand, the contribution from $\theta_{5}$ is determined by the mass spectrum of the additional Higgs bosons and the mixing matrix $O$ as follows;
$d_{e}^{\theta_{5}}\simeq 5.79\times 10^{-29}~{}\mathrm{e\ cm}.$ (107)
Thus, the new physics contribution is given by
$d_{e}\simeq-3\times 10^{-31}~{}\mathrm{e\ cm}.$ (108)
Each contribution is larger than the current constraint $|d_{e}|<4.2\times 10^{-30}~{}\mathrm{e\ cm}$; however the sum of them is lower than it. Consequently, Scenario I avoids the current eEDM constraint although the size of the $CP$-violating phase is not small; $|\theta_{5}|\simeq 0.990$.
We have not fixed some quartic couplings in the Higgs potential because their values are irrelevant in the above discussion. We can choose their value to satisfy the vacuum stability conditions in Eqs. (29)–(31).
V.4.2 Scenario II
In Scenario II, almost all parameters have the same values as in Scenario I. We show only parameters that are different or are not explicitly shown in Scenario I.
The most important difference is the mass of $\eta$. In Scenario II, we use the heavy DM scenario to explain the observed DM relic density, where the relic density is determined by the annihilation into a pair of the additional Higgs bosons. Thus, $\eta$ must be heavier than at least one of the additional Higgs bosons. We assume that
$m_{\eta}=250~{}\mathrm{GeV},$
and the other particles have the same mass as in Scenario I.
$\Lambda_{\alpha}$ is given by
$(\Lambda_{1},\Lambda_{2},\Lambda_{3})\simeq(277,279,284)~{}\mathrm{GeV}.$ (109)
The other parameters required for $h_{i}^{\alpha}$ are assumed to be
$\kappa\tan\beta=30,\quad m_{\nu^{1}}=7.17~{}\mathrm{meV},\quad\delta=1.22\pi,\quad\alpha_{1}=1.07\pi,\quad\alpha_{2}=0.415\pi,$
$\gamma_{1}=1.19\pi,\quad\gamma_{2}=1.97\pi,\quad\gamma_{3}=0.132\pi,\quad\psi_{1}=-0.0852,\quad\psi_{2}=0,\quad\psi_{2}=0.0926\pi.$ (110)
Then the Yukawa couplings $h^{\alpha}_{i}$ are
$\begin{pmatrix}h^{1}_{e}&h^{1}_{\mu}&h^{1}_{\tau}\\ h^{2}_{e}&h^{2}_{\mu}&h^{2}_{\tau}\\ h^{3}_{e}&h^{3}_{\mu}&h^{3}_{\tau}\\ \end{pmatrix}\simeq\begin{pmatrix}2.66e^{0.956\pi i}&0.0137e^{-0.222\pi i}&0.00159e^{0.854\pi i}\\ 2.82e^{0.658\pi i}&0.0473e^{0.755\pi i}&0.00167e^{-0.908\pi i}\\ 2.62e^{-0.102\pi i}&0.0586e^{-0.884\pi i}&0.00207e^{0.971\pi i}\\ \end{pmatrix}.$ (111)
The predictions for the LFV processes are shown in Table. 4, which are consistent with the current upper bounds.
Table 4: Branching ratios for LFV decays in Scenario II
Predictions Upper bounds
$\mu\to e\gamma$ $5.08\times 10^{-14}$ $3.1\times 10^{-13}$ MEGII:2023ltw
$\tau\to e\gamma$ $1.56\times 10^{-15}$ $3.3\times 10^{-8}$ BaBar:2009hkt
$\tau\to\mu\gamma$ $1.33\times 10^{-18}$ $4.4\times 10^{-8}$ BaBar:2009hkt
Predictions Upper bounds
$\mu\to 3e$ $2.79\times 10^{-13}$ $1.0\times 10^{-12}$ SINDRUM:1987nra
$\tau\to 3e$ $1.50\times 10^{-10}$ $2.7\times 10^{-8}$ Hayasaka:2010np
$\tau\to 3\mu$ $7.76\times 10^{-20}$ $2.1\times 10^{-8}$ Hayasaka:2010np
Predictions Upper bounds
$\tau\to e\mu\overline{e}$ $1.60\times 10^{-16}$ $1.8\times 10^{-8}$ Hayasaka:2010np
$\tau\to\mu\mu\overline{e}$ $2.62\times 10^{-17}$ $1.7\times 10^{-8}$ Hayasaka:2010np
$\tau\to ee\overline{\mu}$ $6.95\times 10^{-13}$ $1.5\times 10^{-8}$ Hayasaka:2010np
$\tau\to e\mu\overline{\mu}$ $1.60\times 10^{-16}$ $2.7\times 10^{-8}$ Hayasaka:2010np
The DM relic density can be explained by the freeze-out mechanism via $\eta\eta\to H_{2}H_{2}$, which is induced by $\sigma_{1}$. By taking
$\sigma_{1}=0.0841,$ (112)
we can explain the whole DM relic density by $\eta$. We note that this coupling is almost independent of the constraint from the DM direct detection as discussed in Sec. III.3. The elastic scattering $\eta N\to\eta N$ is induced by $\sigma_{2}$, and the constraint can be avoided by taking the same value as in Scenario I.
Finally, we discuss the eEDM. The contribution from $h_{i}^{\alpha}$ is given by
$d_{e}^{h}=-5.05\times 10^{-29}~{}\mathrm{e\,cm}.$ (113)
$\theta_{5}$ is assumed to be
$\theta_{5}=-0.318\pi\simeq-0.999.$ (114)
The other input parameters for $V_{\mathrm{THDM}}$ are the same as in Scenario I. This change of $\theta_{5}$ leads to a slightly different mass spectrum;
$m_{H_{2}}\simeq 222~{}\mathrm{GeV},\quad m_{H_{3}}=m_{H^{\pm}}\simeq 392~{}\mathrm{GeV}.$ (115)
The eEDM generated by the Higgs sector is evaluated as
$d_{e}^{\theta_{5}}\simeq 4.93\times 10^{-29}~{}\mathrm{e\,cm}.$ (116)
Therefore, the total new contribution is given by
$d_{e}\simeq-1\times 10^{-30}~{}\mathrm{e\,cm},$ (117)
which is consistent with the current limit. Scenario II can also avoid the severe eEDM constraint while keeping large $|\theta_{5}|\simeq 0.999$.
VI Discussions and conclusion
[…]
The LFV processes are also important to test the model. Because of the hierarchy in $h_{i}^{\alpha}$, the processes including $e^{\pm}$ are most significant. In particular, the predicted values for $\mu\to e\gamma$, $\mu\to 3e$, and $\tau\to 3e$ in both benchmark scenarios are close to or larger than the expected sensitivity at MEG-II MEGII:2018kmf , Mu3e Blondel:2013ia ; Mu3e:2020gyw and Belle-II Belle-II:2018jsg , respectively. Thus, we expect that their signal or indication can be detected in each future experiment. We summarize the branching ratio in the benchmark scenarios and the expected sensitivity of these processes in Table 5.
Table 5: The branching ratio for $\mu\to e\gamma$, $\mu\to 3e$, and $\tau\to 3e$ and their expected sensitivity at future planed experiments.
Scenario I Scenario II Expected sensitivity
$\mu\to e\gamma$   $2.95\times 10^{-14}$  $5.08\times 10^{-14}$ $6\times 10^{-14}$ MEGII:2018kmf
$\mu\to 3e$  $4.68\times 10^{-13}$  $2.79\times 10^{-13}$ $1\times 10^{-16}$ Blondel:2013ia ; Mu3e:2020gyw
$\tau\to 3e$  $4.84\times 10^{-10}$  $1.50\times 10^{-10}$ $4\times 10^{-10}$ Belle-II:2018jsg
In this paper, we have investigated two kinds of contributions to the eEDM at the leading order. The size of the next leading order (NLO) contribution is expected to be suppressed by $\mathcal{O}(\alpha)$. In the benchmark scenarios, it is roughly estimated by $d_{e}^{\mathrm{NLO}}=\mathcal{O}(10^{-31})$, which is one order smaller than the current experimental bound. Thus, the interference between $d_{e}^{h}$ and $d_{e}^{\theta_{5}}$ is not largely changed by the NLO contribution. We expect that $\theta_{5}$ can be $\mathcal{O}(1)$ even considering the eEDM at the NLO.
The future ACME-III experiment aims to improve the upper bound on the eEDM to $10^{-30}~{}\mathrm{e\,cm}$. Then, Scenario II can be tested by this improvement. On the other hand, the eEDM in Scenario I is smaller than the expected sensitivity. However, more future experiments would be able to reach the value in Scenario I. In this paper, we have not evaluated the neutron EDM (nEDM) or the proton EDM (pEDM) because they are suppressed by $\cot\beta$ in the quark Yukawa couplings, and the current constraints are weaker than that on the eEDM. Thus, we do not expect that they give an effective constraint on the model currently. However, their future improvements are important to test the $CP$-violation in the model. $h_{i}^{\alpha}$ do not contribute to the nEDM and the pEDM. Therefore, the nEDM and the pEDM can directly access to $\theta_{5}$, which plays an important role in electroweak baryogenesis.
Finally, we shortly comment on EWBG in the model. For successful EWBG, we need to satisfy the second and third conditions of the Sakharov condition by new physics, i.e., large enough $CP$ violation and departure from thermal equilibrium Sakharov:1967dj . $\theta_{5}$ is a source of $CP$ violation in the model. It is $\mathcal{O}(1)$ in the benchmark scenarios and is expected to be large enough to generate the observed baryon asymmetry according to the previous works Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg . The departure from thermal equilibrium is achieved by the strongly first-order EWPT. In the model, EWPT can be strongly first order by the non-decoupling effect of the additional scalar bosons, which also induces the sizable deviation in the Higgs triple coupling Kanemura:2002vm ; Kanemura:2004ch ; Kanemura:2004mg ; Braathen:2019pxr ; Braathen:2019zoh . In our benchmark scenarios, by taking $\mu_{S}=20~{}\mathrm{GeV}$, we have $R_{hhh}\equiv\Delta\lambda_{hhh}/\lambda_{hhh}\simeq 22~{}\%$, where we used the one-loop effective potential to evaluate $\Delta\lambda_{hhh}$. According to Refs. Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg , this would be large enough to make EWPT strongly first-order. Consequently, we expect that EWBG can work in our benchmark scenarios. Since a more detailed study needs complicated analyses using the effective potential, we will study it in our next paper to avoid making this paper too long future .
The non-decoupling effect in $R_{hhh}$ can be tested in the process of di-Higgs production at future high-energy colliders Cepeda:2019klc ; Fujii:2015jha ; Bambade:2019fyw . $R_{hhh}$ in the benchmark scenarios would be tested at the upgraded ILC with the beam energy $\sqrt{s}=1~{}\mathrm{TeV}$, where the expected sensitivity is $10~{}\%$. Strongly first-order EWPT can also be probed by the gravitational waves Grojean:2006bp ; Espinosa:2010hh ; Kakizaki:2015wua ; Hashino:2016rvx ; Hashino:2018wee at future space-based gravitational wave observatories LISA:2017pwj ; Seto:2001qf ; Corbin:2005ny . Furthermore, a sizable non-decoupling effect in EWPT can be tested by the observations of the primordial black holes Hashino:2021qoq ; Hashino:2022tcs
In this paper, the electric dipole moment has been examined in a three-loop neutrino mass model with dark matter originally proposed in Ref. Aoki:2008av . The model contains a $CP$-violating phase in the Higgs potential which plays an important role in electroweak baryogenesis and is thus expected to explain the baryon asymmetry of the Universe simultaneously. However, such a $CP$-violating phase is severely constrained by the measurements of the electron electric dipole moment (eEDM), and a suppression mechanism for the electric dipole moment is necessary to explain the observed baryon asymmetry while avoiding the constraint. In this paper, we examine neutrino mass, lepton-flavor-violating processes, dark matter, and the eEDM in the model. We have shown that the eEDM can be suppressed by destructive interference between the $CP$-violating phases in the Higgs sector and the dark sector with large $CP$-violating phases. We have proposed some benchmark scenarios including $\mathcal{O}(1)$ $CP$-violating phases where tiny neutrino mass and dark matter can be explained while avoiding all the theoretical and experimental constraints. These $CP$-violating phases are expected to be large enough to generate the observed baryon asymmetry in the electroweak baryogenesis scenario.
Acknowledgements
The work of K. E. was supported in part by the National Research Foundation of Korea (Grant No. RS-2024-00352537). The work of S. K. was supported in part by the JSPS KAKENHI Grant No. 20H00160 and No. 23K17691.